### v1

# Deterministic Legal Retrieval: An Action API for Querying the SAT-Graph RAG

Abstract
The Structure-Aware Temporal Graph RAG (SAT-Graph RAG) addresses core limitations of standard Retrieval-Augmented Generation in the legal domain by providing a verifiable knowledge graph that models hierarchical structure, temporal evolution, and causal events of legal norms. However, a critical gap remains: how to reliably query this structured knowledge without sacrificing its deterministic properties. This paper introduces the SAT-Graph API, a formal query execution layer centered on canonical actions—atomic, composable, and auditable primitives that isolate probabilistic discovery from deterministic retrieval. These actions enable: (i) high-precision hybrid search; (ii) robust reference resolution; (iii) point-in-time version retrieval; and (iv) auditable causal tracing. We demonstrate how planner-guided agents can decompose complex queries into Directed Acyclic Graphs (DAGs) of these actions. This two-layer architecture transforms retrieval from an opaque black box to a transparent, auditable process, directly addressing Explainable AI (XAI) requirements for high-stakes domains.
Keywords: Legal AI, Knowledge Graphs, Retrieval-Augmented Generation (RAG), Agentic Frameworks, Explainable AI, Trustworthy AI, Temporal Data Modeling, Legal Ontology.
1 Introduction
The advent of Large Language Models (LLMs) has catalyzed a paradigm shift in information retrieval, with Retrieval-Augmented Generation (RAG) emerging as a dominant architecture for grounding model outputs in factual data [1]. By retrieving relevant text chunks to supplement the model’s parametric knowledge, RAG systems have shown remarkable success in reducing hallucinations. However, the direct application of standard "flat-text" RAG to high-stakes, structured domains such as law reveals fundamental limitations. Legal corpora are not mere collections of text; they are complex, interlinked systems defined by a rigid hierarchical structure, a continuous temporal evolution, and a clear causal chain of legislative actions. Recent studies confirm that traditional RAG frameworks, which rely on surface-level semantic similarity, fail to capture the nuanced relevance required by legal statutes and are insufficient for domain-specific applications without significant adaptation [2, 3].
Standard RAG architectures are ill-equipped to handle this complexity. Even when enhanced with sophisticated semantic segmentation that correctly parses a legal text into its constituent articles and paragraphs, the underlying retrieval mechanism remains fundamentally flawed. It treats these well-defined segments as a simple "flat-text" collection. More critically, it lacks any inherent understanding of versioning. A query about a legal provision is meaningless without a temporal anchor, as the correct answer changes over time. This structural and causal blindness renders such systems fundamentally unreliable for applications demanding legal precision and auditability.
To address these deficiencies, our prior work introduced the Structure-Aware Temporal Graph RAG (SAT-Graph RAG) [4]. This framework moves beyond flat text by modeling legal norms as a verifiable knowledge graph whose ontology explicitly represents the hierarchy, diachronic evolution, and causal events of legislation, as illustrated in Figure 1. While the SAT-Graph provides the necessary verifiable knowledge substrate, a critical gap remains: how can external systems, such as agentic reasoning frameworks or query engines, reliably and flexibly interact with this structured knowledge without sacrificing its deterministic properties? Unstructured access could easily bypass the graph’s guarantees, undermining the very trustworthiness it was designed to provide.
Figure 1: From "flat-text" RAG to SAT Graph RAG
This paper bridges that critical gap. We argue that a trustworthy knowledge base requires a formally defined and auditable interaction protocol. To this end, we introduce a formal query execution layer, which we specify as the SAT-Graph API. The core of this API is a library of canonical actions—atomic, composable, and auditable primitives designed to serve as a secure low-level interface for querying the SAT-Graph. Instead of a monolithic query engine, we provide the essential building blocks that higher-level systems can use to construct complex retrieval plans. These actions provide a comprehensive toolkit for: (i) high-precision hybrid search combining semantic, lexical, and structured metadata filters; (ii) robust resolution of textual references to formal identifiers; (iii) deterministic retrieval of point-in-time versions; and (iv) complete, auditable tracing of causal lineage.
We demonstrate how a planner-guided agent, such as those proposed in advanced Graph RAG frameworks [5], can leverage this API by decomposing a user’s natural language query into an executable Directed Acyclic Graph (DAG) of these actions. This two-layer architecture—with the SAT-Graph as the verifiable knowledge base and our Action API as the secure interaction protocol—establishes a new blueprint for building trustworthy legal AI. It is a direct architectural response to the broader challenge of Explainable AI (XAI), moving the system’s reasoning from an opaque "black box" to a transparent process that can be audited and understood, a critical requirement for high-stakes domains [6]. This allows advanced reasoning engines to operate on legal knowledge while ensuring every step of the retrieval process remains explicit, verifiable, and grounded in the structure of the law.
The main contributions of this paper are as follows:
  • We formally specify a Canonical Action API, a library of primitive and composable actions for interacting with legal knowledge graphs, covering primitives for retrieval, temporal resolution, and causal analysis.
  • We demonstrate how this API serves as a foundational layer for agentic architectures, enabling them to decompose complex questions into verifiable retrieval plans.
  • We articulate how this architecture decouples the knowledge representation (the graph) from the query logic (the agent), enabling a robust and explainable approach to answering legal questions with LLMs.
The remainder of this paper is structured as follows. Section 2 discusses related work in legal AI and graph-based retrieval. Section 3 details the specification of our Canonical Action API. Section 4 presents use cases demonstrating its application, and Section 5 concludes with a discussion of the implications and future directions.
2 Related Work
Our work is situated at the intersection of legal informatics, retrieval-augmented generation, knowledge graphs, and agentic AI systems. We position our contribution by reviewing the state of the art in these areas.
2.1 AI in Legal Information Retrieval
The application of artificial intelligence to legal text is a long-standing field of research. Early systems relied on handcrafted rules and expert knowledge to perform legal reasoning [7]. While the advent of machine learning brought significant advances through models like Legal-BERT [8], a common thread remains the treatment of legal documents as unstructured text. This perspective often overlooks the deep, formal semantics embedded in the law’s structure.
Our approach, in contrast, aligns with a mature research tradition that seeks to formalize legal knowledge using ontologies. The challenge of managing the "diachronic evolution" of legal norms has long been recognized, with early work proposing ontologies to identify and track legal resources over time [9]. Our use of a formal, LRMoo-based model [10] builds upon this foundation. Furthermore, the explicit modeling of legislative changes via our ‘Action‘ nodes is a direct response to the recognized need for formal languages to describe and manage changes in complex knowledge graphs, a problem addressed by frameworks like the Knowledge Graph Change Language (KGCL) [11]. By adopting a formal, top-down ontological model, we enable more sophisticated, machine-readable applications that respect the inherent structure of the law.
2.2 Retrieval-Augmented Generation and its Limitations
Retrieval-Augmented Generation (RAG) has become the de-facto standard for enabling LLMs to answer questions over specific corpora [1]. In its canonical form, RAG operates by embedding and indexing text chunks, retrieving the most relevant via vector similarity search.
While effective for unstructured text, this "flat-text" RAG paradigm falters in the legal domain. Recent benchmarks confirm that the retrieval stage remains a critical bottleneck for legal RAG systems [12]. Our claims of "context blindness" and "temporal naivety" are not hypothetical; they reflect well-documented failure modes. Studies show that a reliance on surface-level similarity fails to capture the "nuanced relevance" of legal statutes [3] and that standard RAG is insufficient without domain-specific adaptations to handle the complexity of legal queries [2]. The inability of these systems to guarantee the retrieval of the correct historical version of a law is a fatal flaw for any serious legal application.
2.3 Knowledge Graphs for Retrieval (Graph RAG)
To overcome the limitations of flat-text retrieval, a growing body of work has focused on Graph RAG [13]. Frameworks like Youtu-GraphRAG have demonstrated the power of this approach [5]. The architectural advantage of a knowledge graph is that it provides a persistent, structured, and connected memory that an external agent can query, enabling more intricate reasoning tasks than are possible with isolated text chunks [14].
Our SAT-Graph RAG [4] aligns with this trend but makes a critical distinction. Unlike generalist approaches that discover a graph "bottom-up" from unstructured text, the SAT-Graph is constructed "top-down," mapping the pre-existing, formal structure of the law onto a deterministic, temporal ontology. The most important entities are not named persons or places, but the structural components of the law itself and the legislative events that alter them. While this provides a robust knowledge base, the question of a standardized interaction protocol remains open.
2.4 Agentic Frameworks for Reasoning and Retrieval
Recent advances have seen LLMs employed not just as generators, but as reasoning agents capable of planning and tool use [15]. These agents can decompose complex questions and utilize a "library of tools" to gather information. The API we propose is, in effect, this specialized tool library for the legal graph.
A critical insight from recent research is the symbiotic relationship between architectural innovation and evaluation benchmarks. The legal domain has historically suffered from a lack of benchmarks that require deep, multi-hop reasoning over structured, temporal data. However, new, more challenging benchmarks are emerging, such as "Bar Exam QA" and "Housing Statute QA," which are specifically designed to require "reasoning-intensive" queries rather than simple lexical overlap [16]. These new benchmarks expose the deep limitations of traditional RAG and create a clear need for architectures like ours. The SAT-Graph, combined with our Canonical Action API, is precisely the type of verifiable, structure-aware framework needed to successfully address these next-generation evaluation tasks.
2.5 Positioning Our Contribution
Our work synthesizes these threads. We build upon the foundation of our SAT-Graph, which addresses the representation problem in legal AI. We acknowledge the power of agentic reasoning frameworks for flexible query resolution. However, we identify a critical missing component: the secure and auditable bridge between the reasoning agent and the structured knowledge.
While prior work has focused on either the graph structure itself or the high-level agentic logic, our contribution is the formal specification of the intermediate layer: a Canonical Action API. This is not another monolithic retrieval system, but a low-level protocol that provides the necessary safety guarantees for the legal domain. It enables advanced agents to perform complex reasoning while ensuring that every interaction with the knowledge base is explicit, verifiable, and grounded in the temporal and structural reality of the law.
3 Specification of the Canonical Action API
The central thesis of our work is that trustworthy interaction with a structured knowledge base like the SAT-Graph requires a formal, auditable protocol. We introduce this protocol as the Canonical Action API, a library of atomic and composable actions. This API serves as an abstraction layer, hiding the underlying graph database implementation (e.g., Cypher queries over a Neo4j database). Crucially, it provides a unified interface for both probabilistic discovery actions (like resolving natural language references) and fully deterministic retrieval actions that operate on formal identifiers. In this pattern, an external agentic framework is designed to interact with the knowledge base primarily through these actions, ensuring that every step in a retrieval plan is explicit, verifiable, and its level of certainty is understood, as shown in Figure 2.
Figure 2: Diagram illustrating a reasoning agent decomposing a user prompt into tasks that are executed through actions provided by the SAT-Graph API.
3.1 Design Principles
The design of the API is guided by three core principles that, together, ensure trustworthiness even when interfacing with probabilistic language models.
  • Maximal Determinism: We acknowledge that any system interpreting natural language has inherently probabilistic components. Our design philosophy is to isolate this non-determinism at the API’s entry points. Actions for discovery and search are probabilistic by nature. However, the core architectural contribution of this API is to enable the orchestration of fully deterministic retrieval plans once a formal identifier is obtained. It is this capability for achieving determinism in the final, grounded query that fundamentally distinguishes our approach and justifies the "Deterministic Retrieval" paradigm. Once a formal id (URI/URN) is acquired, all subsequent actions that operate on it are guaranteed to be deterministic.
  • Composability: Actions are designed as atomic "building blocks." They can be chained together by a planning agent to construct complex query workflows. This allows for flexibility while ensuring that the overall plan is composed of individually verifiable steps, clearly distinguishing between initial probabilistic grounding and subsequent deterministic traversal. Server-side execution of composed actions ensures efficient processing even for complex multi-step plans.
  • Verifiability through Auditability: This principle is our safeguard against the initial non-determinism. Every action, including the probabilistic ones, must return not just a result but also a confidence score or justification. When an agent executes a plan, the sequence of action calls and their outputs forms a complete, human-readable audit trail. This log makes the entire process verifiable, even if not every step is deterministic. A user can inspect the log and see, "The system resolved ’Article 6’ to id ’…’ with 98% confidence," allowing them to trust or question the initial grounding step.
3.2 Core Data Models
The actions in our API operate on and return a set of core data objects that correspond directly to the entities in the SAT-Graph ontology, as illustrated in Figure 3. The ontology makes a clear distinction between versioned normative entities (Item) and atemporal conceptual entities (Theme).
Figure 3: SAT-Graph Ontology
Item:
Represents a versioned, structural entity of a legal norm. It is instantiated as one of two concrete subtypes: Work or Work Component.
  • id (ID): A unique, canonical identifier (e.g. URI/URN).
  • category (string): The architectural class (‘Work‘ or ‘Work Component‘).
  • type (string): The specific semantic subtype (‘Constitution‘, ‘Article‘, ‘Chapter‘).
  • label (string): The human-readable label.
  • parent (id)?: (Optional) The id of its single structural parent. A ‘Work‘ has no parent.
  • children (list[id])?: (Optional) A list of ids for its structural children. Leaf nodes have no children.
  • metadata (JSON)?: (Optional) A flexible json store for structured, atemporal properties of the item—examples of metadata description of a legal norm’s Work in [17].
Theme:
Represents an atemporal, conceptual entity used to classify Items and to organize knowledge. Themes can be arranged in their own taxonomy.
  • id (ID): A unique, canonical identifier.
  • label (string): The human-readable label.
  • parent (list[id])?: (Optional) A list of ids of broader themes. Themes can have multiple parents, creating a directed acyclic graph (DAG) structure. Root themes have an empty list.
  • children (list[id])?: (Optional) A list of ids of more specific themes.
  • members (list[id])?: (Optional) A list of ids of Items classified under this theme.
Architectural Note: Separation of Normative and Conceptual Entities
A foundational design principle of this framework is the strict separation of normative entities (Item) from conceptual, classificatory entities (Theme). Items represent the formal, versioned structure of the law itself. Themes, in contrast, form an independent, atemporal ontology used to organize and discover knowledge. An Item’s structural relationships (‘parent‘, ‘children‘) define its fixed position within a document, forming a strict hierarchy. A Theme’s relationships define both its taxonomic position relative to other themes (‘parent‘, ‘children‘) and its classificatory relationship to the law (‘members‘). This clear separation allows an agent to navigate the formal structure of a norm (or component) and the conceptual map of the legal domain through distinct, unambiguous API actions.
The ‘Version‘ (Item Temporal Version) and ‘Action‘ data models build upon this DAG foundation to create a fully versioned and causally-linked graph.
Version:
Represents a specific, time-bound version of an Item. An Version captures not only the content but also the structural position of an item at a specific point in time.
  • id (ID): A unique identifier for this specific version.
  • item_id (ID): The id of the abstract Item it is a version of.
  • validity_interval ([date, date])?: A tuple representing the start and end dates of its validity. The end date is optional (or null) to represent a currently valid or open-ended version.
  • parent (list[ID])?: (Optional) A list of identifiers for its parent Versions. A ‘Work‘’s Version has no parents.
  • children (list[ID])?: (Optional) A list of identifiers for its children Versions, representing the complete structure at this point in time.
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the version.
Action:
Represents a reified legislative event as the most granular, atomic unit of change in the graph. It is a first-class entity that models a single, auditable causal link, connecting a specific clause of an authorizing law to the specific change it enacts.
  • id (ID): A unique identifier for the action.
  • type (string): The type of legislative action (e.g., ‘Amendment‘, ‘Revocation‘, ‘Creation‘).
  • date (date): The effective date of the action.
  • source_version_id (ID): The identifier for the specific Version of the legal text that authorizes this single change.
  • terminates_version? (ID): (Optional) The specific Version whose validity is terminated by this action.
  • produces_version (ID): The specific Version that is created by this action.
TextUnit:
Represents a concrete piece of textual information. It is a flexible object used to hold any "aspect" of text associated with any primary entity in the graph. This could be the canonical text of a versioned item (an Version), a detailed description of a Theme, or a textualized metadata aspect of an Item (like a publication info.).
  • id (string): A unique identifier for this specific piece of text.
  • source_node_type (string): The type of the source node (e.g., "Item", "Version", "Action", "Theme"), to allow for unambiguous retrieval by an agent.
  • source_node_id (ID): The unique identifier of the node in the graph to which this text is attached.
  • language (string): The language of the content (e.g., "pt-br").
  • aspect (string): The semantic role of the text relative to its source node (e.g., "canonical" text of a Version, "description" of a Theme, "textual_metadata" of a Work).
  • content (string): The raw textual content.
3.3 API Action Specification
The API is organized into functional categories based on agentic intent. A core design guarantee governs all actions: any action that accepts a formal ID as its primary input is fully deterministic. The only exceptions are the discovery and search actions, which interpret natural language text. This section details the signature and purpose of each action.
Bridging Baseline RAG with a Temporal Knowledge Graph
A key design challenge is to provide a search interface that is both simple enough to replicate the behavior of a standard, "temporally-naïve" RAG system, while also exposing the full power of the underlying temporal graph. Traditional RAG pipelines typically index a single, current version of a text. A query like “Which social rights are guaranteed?” executes a semantic search over these current texts.
The SAT-Graph, however, contains all historical versions. A simple semantic search would return a mix of current and outdated, yet relevant, texts. To manage this complexity, the API’s primary search primitive, searchTextUnits, is explicitly designed to handle two core user intents through a simple timestamp parameter:
  • Default (Current State Query): When the timestamp is omitted, the search is automatically constrained to only the currently valid texts, thus replicating the expected behavior of a baseline RAG.
  • Temporal Query: When a timestamp is provided, the search is constrained to the historical snapshot of texts that were valid on that specific date.
This design provides a single, powerful action that elegantly bridges the gap between simple RAG queries and complex historical analysis.
3.3.1 Discovery and Search Actions
These actions serve as the primary entry points for an agent, used when it needs to find entities based on ambiguous natural language references or semantic content.
  1. 1.
    resolveItemReference(reference_text: string, context_id?: ID,
    top_k?: int) $\rightarrow$ list[{ID, confidence: float}]
    • Description: Translates an ambiguous, natural-language reference to a document/component into a ranked list of candidate Item ids. This is the critical first step for any query that names a specific item, acting as the bridge between user language and the graph’s formal identifiers.
    • Parameters:
      • reference_text: A string such as "Article 5 of the Constitution", "the Civil Code", or a relative reference like "the previous section".
      • context_id: (Optional) The id of an Item that provides a structural context. This allows the system to accurately resolve relative references and disambiguate ambiguous ones (e.g., resolving "Article 10" to the one within the same law as the context id).
      • top_k: (Optional) The maximum number of candidate matches to return (e.g., defaults to 3).
    • Returns: A ranked list of objects, each containing a probable id and a confidence score indicating the certainty of the match.
    • Discussion: This action is indispensable for grounding agentic workflows. Document/component references are often highly ambiguous ("Article 1" could exist in hundreds of laws) or relative ("the following paragraph"). This function handles this by returning a ranked list of candidates rather than forcing a single, potentially incorrect, guess. The optional context_id parameter significantly enhances its precision, enabling sophisticated, context-aware interactions. This design makes the entire system safer and more robust, preventing critical errors that arise from incorrect initial grounding.
  2. 2.
    resolveThemeReference(reference_text: string, top_k?: int) $\rightarrow$
    list[{ID, confidence: float}]
    • Description: Translates a natural-language name or label for a legal theme into a ranked list of candidate Theme ids. This is the primary entry point for grounding a query in a specific conceptual category.
    • Parameters:
      • reference_text: A string representing the name of a theme, such as "Social Security" or "Environmental Law".
      • top_k: (Optional) The maximum number of candidate matches to return (e.g., defaults to 3).
    • Returns: A ranked list of objects, each containing a probable Theme id and a confidence score.
    • Discussion: This action is the conceptual counterpart to resolveItemReference. It performs entity linking for thematic nodes, allowing an agent to deterministically identify a theme mentioned by name before using it as a scope in other actions like searchTextUnits.
  3. 3.
    searchTextUnits(version_ids?: list[ID], item_ids?: list[ID],
    theme_ids?: list[ID], metadata_filter?: MetadataFilter,
    timestamp?: date, semantic_query?: string, lexical_query?: string,
    language?: string, aspects?: list[string], top_k?: int) $\rightarrow$
    list[{text_unit: TextUnit, score: float}]
    • Description: Performs a hybrid search to discover TextUnits nodes by combining semantic, lexical, and structured criteria returning a ranked list of the most relevant results, ready to be used as context for a generation model. This is the primary, optimized entry point for all standard RAG queries.
    • Parameters:
      • version_ids: (Optional) A list of specific Version (Item Temporal Version) identifiers. Provides an explicit, pre-resolved versioned scope. This parameter is mutually exclusive with the combination of item_ids/theme_ids and timestamp.
      • item_ids: (Optional) A list of Item ids. Defines the structural scope. The search includes these items and their descendants.
      • theme_ids: (Optional) A list of Theme ids. Expands the search scope to include all member Items of the specified themes and all of their descendant themes. For example, searching within the "Tribute Law" theme will automatically include items from its sub-themes like "Federal Taxes" and "State Taxes".
      • metadata_filter: (Optional) A structured object (MetadataFilter) that specifies structured filters to be applied. This object can contain clauses like:
          item_metadata_filter (dict): (Optional) A set of key-value pairs to filter against the metadata of the Items (e.g., ‘"jurisdiction": "federal"‘).
          version_metadata_filter (dict): (Optional) A set of key-value pairs to filter against the metadata of the Version itself (e.g., ‘"publication_date": ">=2010-01-01"‘).
      • timestamp: (Optional) A date to constrain the search when using item_ids or theme _ids. If omitted, defaults to the current date ("now"). This parameter is ignored if version_ids is provided.
      • semantic_query (string): (Optional) The natural language query string.
      • lexical_query (string): (Optional) A set of keywords or phrase for a full-text syntactic search.
      • language: (Optional) A code for the desired language.
      • aspects: (Optional) A list of strings specifying the desired textual aspects (e.g., [‘"canonical"‘, ‘"summary"‘]). Defaults to [‘"canonical"‘].
      • top_k: (Optional) The maximum number of results to return.
    • Returns: A ranked list of objects, each containing a matching TextUnit and its similarity score.
    • Discussion: This action’s dual-mode design for scoping provides maximum flexibility.
      When to use Mode 1 (Explicit Version Scope): Use when you have already identified a specific set of Versions (e.g., from ‘getItemHistory‘) and need to search exclusively within that curated set. This provides a direct, high-performance path for advanced analytical tasks.
      When to use Mode 2 (Conceptual Scope): Use for natural language queries like "What did Law X say about topic Y in 2005?" The agent defines the "what" (item_ids/theme_ids) and the "when" (timestamp), and the API handles temporal resolution automatically.fic set of Versions of interest (e.g., from getItem History) and needs to perform a semantic search exclusively within that curated set. Using this mode provides a direct, high-performance path for advanced analytical tasks.
      Furthermore, the aspects parameter enables a truly holistic search across the entity hierarchy. By default, the search targets the "canonical" text of the relevant Versions. However, by specifying other aspects (e.g., ["canonical", "textual_metadata"]), the search space expands to include all TextUnits matching those aspects, regardless of whether they are attached to the filtered Versions themselves or to their timeless parent Items. This allows an agent to, in a single query, find a match in a Work’s "publication info." just as easily as in an Version’s official text, creating a comprehensive and context-aware retrieval process.
  4. 4.
    searchItems(item_ids?: list[ID], theme_ids?: list[ID],
    item_metadata_filter?: dict, semantic_query?: string,
    lexical_query?: string, top_k?: int) $\rightarrow$
    list[{item: Item, score: float}]
    • Description: Performs a hybrid search to discover Item nodes (the structural entities of the graph) by combining semantic, lexical, and structured criteria. This action searches across the entire version history of each item’s text to ensure comprehensive discovery.
    • Parameters:
      • item_ids: (Optional) A list of Item ids to define the search scope.
      • theme_ids: (Optional) A list of Theme ids to expand the search scope.
      • item_metadata_filter (dict): (Optional) A set of key-value pairs to filter against the metadata of the Item (e.g., ‘"jurisdiction": "federal"‘).
      • semantic_query (string): (Optional) The natural language query string.
      • lexical_query (string): (Optional) A set of keywords or phrase for a full-text syntactic search.
      • top_k: (Optional) The maximum number of results to return.
    • Returns: A ranked list of objects, each containing a matching Item and its similarity score.
    • Discussion: This action’s purpose is structural discovery, not content retrieval. When an agent uses searchItems, it is attempting to find the locus of a concept within the graph’s formal structure. To support this, the search must be comprehensive, and is therefore intentionally designed to be time-agnostic, considering content from an item’s entire version history.
      The distinction from searchTextUnits is critical:
      • To find the text about "housing" that was valid in 2005, an agent must use searchTextUnits( ..., timestamp="2005-01-01").
      • To find the item that, at any point in its history, has dealt with "housing", an agent must use searchItems.
      Adding a timestamp to this function would create a confusing overlap and undermine its core purpose as a high-recall structural discovery tool. The item_metadata_filter parameter, in contrast, enhances this purpose by enabling high-precision discovery based on the atemporal metadata of the structural nodes themselves.
  5. 5.
    searchThemes(semantic_query?: string, lexical_query?: string,
    top_k?: int) $\rightarrow$ list[{theme: Theme, score: float}]
    • Description: Performs a hybrid search to discover Theme nodes based on a descriptive query. This action searches exclusively on the textual descriptions associated with the themes themselves, not on the content of their member items.
    • Parameters:
      • semantic_query (string): (Optional) The natural language query string.
      • lexical_query (string): (Optional) A set of keywords or phrase for a full-text syntactic search.
      • top_k: (Optional) The maximum number of candidate themes to return.
    • Returns: A ranked list of candidate objects, each containing a theme’s id and a confidence score.
    • Discussion: This action’s purpose is to map a user’s description of a topic to the most appropriate formal concept in the graph’s ontology. To ensure a precise and high-performance conceptual match, the search is intentionally scoped to the curated textual descriptions of the Theme nodes. Searching the content of all member Items would introduce significant semantic noise and lead to inaccurate results, as large themes could overshadow more specific ones. This design choice maintains a clear separation of concerns: use searchThemes to find the right conceptual category, then use searchTextUnits with the resulting theme’s id to find relevant content within that category.
3.3.2 Deterministic Fetch Actions
Once an agent has acquired a specific ID, these actions are used to retrieve the full, corresponding data objects in a deterministic, non-ambiguous way.
  1. 6.
    getItem(id: ID) $\rightarrow$ Item
    • Description: Retrieves the full Item object given its unique id.
    • Parameters:
      • id: The canonical id of the item.
  2. 7.
    getTheme(id: ID) $\rightarrow$ Theme
    • Description: Retrieves the full Theme object given its unique id.
    • Parameters:
      • id: The canonical id of the theme.
  3. 8.
    getAction(id: ID) $\rightarrow$ Action
    • Description: Retrieves the full Action data object given its unique identifier. This provides access to all details of a action, including its precise ‘source_version_id‘.
    • Parameters:
      • id: The unique identifier of the action.
    • Returns: The corresponding Action data object.
  4. 9.
    getValidVersion(item_id: ID, timestamp: date-time, policy:
    TemporalPolicy) $\rightarrow$ Version
    • Description: The core temporal resolution function. It finds the single Version that was legally in force on a specific date.
    • Parameters:
      • item_id: The id of the item of interest.
      • timestamp: The specific date for which to check validity.
      • policy: An enum specifying the resolution logic, e.g., ‘TemporalPolicy.SnapshotLast‘ which finds the version whose validity interval contains the timestamp.
    • Returns: The unique Version object valid on the given date.
    • Discussion: The explicit ‘policy‘ parameter is essential for guaranteeing determinism. It allows the agent to specify the exact temporal logic required, such as retrieving the last valid version within a day (snapshot semantics) or handling boundary conditions in a predictable manner. The policy parameter controls temporal resolution: PointInTime: Returns the version valid at exactly the specified timestamp. SnapshotLast: Returns the last known version before or at the timestamp. Default: PointInTime.
  5. 10.
    getTextForVersion(version_id: ID, language: string) $\rightarrow$ TextUnit
    • Description: Deterministically retrieves a specific textual representation (a TextUnit) for a single, known Version. This is a direct-access, primary-key-based lookup, not a search operation.
    • Parameters:
      • version_id: The unique identifier of the Version object whose text is requested.
      • language: A code for the desired language (e.g., "pt-br", "en").
    • Returns: The single, corresponding TextUnit object.
    • Discussion: This action is fundamentally different from searchTextUnits. While the latter is a probabilistic discovery tool based on semantic similarity, getTextForVersion is a deterministic fetch operation. It serves the critical use case where an agent has already identified a specific Version (e.g., via getValidVersion or getItemHistory) and requires its exact textual content. Maintaining this as a distinct, highly-optimized action ensures API clarity, guarantees determinism, and provides a clear audit trail of the agent’s execution plan.
3.3.3 Structural Navigation Actions
These actions are used to traverse the atemporal, hierarchical, and thematic structures of the graph.
  1. 11.
    enumerateItems(item_ids?: list[ID], theme_ids?: list[ID],
    depth?: int) $\rightarrow$ list[Item]
    • Description: Enumerates the structural or thematic members of a given scope. This is the primary action for navigating the graph’s hierarchy and membership relationships.
    • Parameters:
      • item_ids: (Optional) A list of Item ids. Defines the structural scope. The search includes these items and their descendants.
      • theme_ids: (Optional) A list of Theme ids. Expands the scope to include all member items of these themes.
      • depth: (Optional) An integer controlling the traversal depth for hierarchical Item scopes. If ‘depth=1‘, only direct children are returned. If omitted or negative, all descendants are returned recursively. This parameter is ignored for Theme scopes.
    • Returns: A list of Item objects. The behavior is polymorphic based on the type of the input ‘scope_id‘:
      • For a structural Item scope: it returns its hierarchical children/descendants.
      • For a Theme scope: it returns all of its member Items.
    • Discussion: This action’s purpose is structural enumeration and navigation, not search filtering. It is the correct tool for tasks like building a table of contents or programmatically exploring the members of a conceptual theme. For performing content-based searches within a scope, agents should pass the scope’s id directly to the item_ids or theme_ids parameters of the search functions (searchTextUnits, searchItems), as this is significantly more performant than retrieving the members first and then passing a long list of IDs.
  2. 12.
    getAncestors(item_id: ID) $\rightarrow$ list[Item]
    • Description: Navigates the hierarchy upwards from a specific structural Item to its parent, grandparent, and so on, up to the root Work. This is crucial for retrieving the full contextual path of any given provision.
    • Parameters:
      • item_id: The id of the starting Item.
    • Returns: An ordered list of Item objects representing the hierarchical path, from the highest-level ancestor (e.g., a ‘Title‘) down to the immediate parent of the input item. The root Work itself is typically excluded.
    • Discussion: This action is the fundamental tool for contextualization. The legal meaning of a provision is often determined by the Chapter or Title it belongs to. After a search action like searchTextUnits identifies a relevant item, an agent should use getAncestors to retrieve its structural "breadcrumb". This allows the generation of richer, context-aware responses, significantly improving the quality and explainability of the final answer.
  3. 13.
    getThemesForItem(item_id: ID) $\rightarrow$ list[Theme]
    • Description: Retrieves a list of all Themes that are directly associated with a specific Item. This provides the inverse navigation from the structural hierarchy to the conceptual ontology.
    • Parameters:
      • item_id: The id of the Item whose thematic classifications are requested.
    • Returns: A list of all Theme objects that classify the given Item.
    • Discussion: This action is the essential counterpart to using a Theme as a scope in enumerate Items. It makes the relationship between items and themes fully bidirectional, enabling an agent to efficiently answer questions like "What topics does this article cover?". Without this primitive, an agent would need to resort to a highly inefficient brute-force search across the entire thematic ontology. This function completes the API’s structural navigation toolkit.
  4. 14.
    getItemContext(item_id: ID) $\rightarrow$ StructuralContext
    • Description: Retrieves the complete, immediate structural context of a single Item in one efficient call. This includes its parent, siblings, and direct children.
    • Parameters:
      • item_id: The id of the Item for which to retrieve the context.
    • Returns: A StructuralContext object containing the full data for the target item and its immediate hierarchical neighbors.
      • target (Item): The full object of the item in question.
      • parent (Item)?: The full object of its immediate parent.
      • siblings (list[Item])?: A list of full objects for its siblings (other items with the same parent).
      • children (list[Item])?: A list of full objects for its immediate children.
    • Discussion: This is a high-utility, performance-oriented action designed for applications that need to render a contextual view of an item (e.g., a document browser or a table of contents). Replicating this functionality would require multiple, less efficient API calls to fetch ancestors, descendants, and then compute the sibling set. By providing this as a single, atomic operation, the API dramatically simplifies the logic for common UI and agentic navigation tasks.
3.3.4 Causal & Lineage Analysis Actions
These actions enable agents to traverse the temporal and causal dimensions of the SAT-Graph, understanding the history and evolution of legal items.
  1. 15.
    getItemHistory(item_id: ID) $\rightarrow$ list[Action]
    • Description: Provides a complete, ordered timeline of all actions that have affected a given Item.
    • Parameters:
      • item_id: The id of the Item whose history is requested.
    • Returns: A time-ordered list of all Action objects that have either created or terminated a version (Version) of the specified Item.
    • Discussion: This action provides a complete, auditable narrative of an item’s evolution. An agent can iterate through the returned list of Actions to reconstruct every state transition, accessing the "before" state via ‘action.terminates_version‘ and the "after" state via ‘action.produces_version‘ for each step.
  2. 16.
    traceCausality(version_id: ID) $\rightarrow$ {creating_action: Action,
    terminating_action?: Action}
    • Description: Traces the direct causal links for a single Version. It identifies the specific Action that brought this Version into existence (its provenance) and, if applicable, the Action that terminated its validity.
    • Parameters:
      • version_id: The unique identifier of the Version.
    • Returns: An object containing the Action that created the Version and, optionally, the Action that terminated it.
    • Discussion: This action is crucial for pinpointing the exact legislative act responsible for a specific version of a legal text, facilitating precise auditability.
  3. 17.
    getVersionsInInterval(item_id: ID, start_date: date,
    end_date: date) $\rightarrow$ list[Version]
    • Description: Retrieves all Versions of a given Item that were legally valid at any point within a specified time interval.
    • Parameters:
      • item_id: The id of the item of interest.
      • start_date: The start date of the time interval for the query.
      • end_date: The end date of the time interval for the query.
    • Returns: A time-ordered list of all Version objects whose validity period overlaps with the specified interval.
    • Discussion: This function serves as a critical performance optimization for historical analysis. While getItemHistory retrieves the causal events (‘Actions‘), this action retrieves the resulting states (‘Versions‘) within a specific temporal window. Replicating this on the client-side would require fetching the entire history and performing complex filtering. By providing this as a server-side primitive, the API enables an agent to efficiently answer questions like "show all versions of this article during a specific presidential term," which is essential for building responsive and scalable legal research applications.
  4. 18.
    compareVersions(version_id_A: ID, version_id_B: ID) $\rightarrow$ TextDiffReport
    • Description: Performs a granular, structure-aware comparison between the textual content of two different Versions of the same Item. It highlights additions, deletions, and modifications.
    • Parameters:
      • version_id_A: Identifier of the first Version.
      • version_id_B: Identifier of the second Version.
    • Returns: A structured TextDiffReport object detailing the changes.
    • Discussion: This action provides a powerful, server-side analysis that goes beyond a simple textual ‘diff‘. Because the server has access to the full graph structure, it can perform a structural comparison, identifying not just changed words but also added or removed components (e.g., a new paragraph). This transforms the comparison from a mere text-level operation into a true structural impact analysis, which is fundamental for legal practitioners. Centralizing this complex logic ensures high performance and a consistent, high-quality analysis for all client applications.
  5. 19.
    getActionsBySource(source_work_id: ID, action_types?: list[string]) $\rightarrow$ list[Action]
    • Description: An essential action for impact analysis that provides forward causal tracing. Given the id of a source Work (e.g., an amending law), it retrieves all atomic Actions authorized by any clause within that Work.
    • Parameters:
      • source_work_id: The id of the source Work (e.g., "Law No. 50 of 2001").
      • action_types: (Optional) A list of strings to filter the results to specific types of actions (e.g., [‘Revocation‘, ‘Creation‘]).
    • Returns: A comprehensive list of all Action objects caused by the specified source that match the filter criteria.
    • Discussion: This action is the logical inverse of traceCausality. While the latter traces backwards from a single effect to its granular cause, getActionsBySource traces forwards from a high-level cause (an entire legislative document) to all of its granular effects. It is the primary, high-performance primitive for answering the critical question: "What did Law X do?". The optional action_types filter allows for more targeted impact analysis, such as isolating only the revocations or creations enacted by a specific law.
3.3.5 Macro-Level Analysis Actions
Beyond entity retrieval and causal tracing, the Canonical Action API provides a distinct class of actions designed for high-level, aggregate analysis. These functions represent a deliberate architectural choice to offload computationally intensive processing from the client to the server. While an agent could theoretically replicate their logic by making numerous, sequential calls to lower-level primitives, doing so would be highly inefficient and impractical for large-scale analysis.
By treating server-side aggregation as a core primitive, the API enables agents and applications to efficiently answer broad, macro-level questions about legislative evolution. Actions in this category typically do not return lists of full entity objects, but rather structured, lightweight summary objects (like an ImpactReport) containing aggregate statistics and lists of identifiers, designed to be selectively hydrated by the agent as needed.
  1. 20.
    summarizeImpact(item_ids?: list[ID], theme_ids?: list[ID],
    time_interval: TimeInterval, action_types?: list[string]) $\rightarrow$
    ImpactReport
    • Description: An aggregate analysis action that provides a high-level summary of changes that occurred within a specific scope during a defined time period.
    • Parameters:
      • item_ids: (Optional) A list of Item ids to define the structural scope of the analysis.
      • theme_ids: (Optional) A list of Theme ids to define the thematic scope of the analysis.
      • time_interval: A start and end date for the analysis window.
      • action_types: (Optional) A list of strings to filter the summary to specific action types (e.g., [‘Revocation‘, ‘Creation‘]). If omitted, all action types are included.
    • Returns: A structured ImpactReport object, containing aggregate statistics (e.g., counts of each action type) and lists of identifiers for the relevant Actions and affected Items.
    • Discussion: This powerful action enables macro-level analysis and is a deliberate design choice to offload complex processing to the server. Replicating its logic on the client-side would require numerous, inefficient calls. By treating server-side aggregation as a core primitive, the API ensures efficiency and simplifies agent logic for this common legislative analysis use case. The ImpactReport intentionally contains lightweight identifiers, allowing the agent to selectively hydrate full details using batch operations like getBatchActions as needed.
3.3.6 Introspection & Metadata Actions
To enable agents to be more adaptive and robust, the API provides a set of introspection actions that allow them to discover the capabilities and boundaries of the knowledge graph at runtime.
  1. 21.
    getTemporalCoverage(item_id: ID) $\rightarrow$ TimeInterval
    • Description: Returns the start and (optional) end dates of the known version history for a specific Item.
    • Returns: A tuple containing the start date of the first Version and the end date of the last known Version. The end date is null if the item is still considered active.
    • Discussion: This action enables an agent to be aware of its own knowledge boundaries. It should be used to proactively validate a user’s temporal query before attempting retrieval, allowing for graceful handling of requests for dates outside the known history and improving the overall user experience.
  2. 22.
    getAvailableLanguages() $\rightarrow$ list[string]
    • Description: Returns a list of all language codes (e.g., "pt-br", "en") for which textual content is available in the graph.
    • Discussion: This enables an agent or user interface to dynamically present language options to the user, ensuring that requests are only made for languages that the system actually supports.
  3. 23.
    getSupportedActionTypes() $\rightarrow$ list[string]
    • Description: Returns a list of all action types (e.g., ‘Amendment‘, ‘Revocation‘) recognized by the graph’s ontology.
    • Discussion: This action provides the canonical list of values that can be used in the action_types filter of other functions like summarizeImpact. It allows a client application to dynamically build filtering interfaces without hardcoding ontology-specific values.
  4. 24.
    getRootThemes() $\rightarrow$ list[Theme]
    • Description: Retrieves a list of the full Theme objects that are at the root of the thematic taxonomy (i.e., themes that have no parents).
    • Returns: A list of all root Theme objects.
    • Discussion: This action serves as the primary and most efficient entry point for an agent or user interface to explore the thematic ontology. By returning the full root objects, it allows a client to render the top-level categories without requiring subsequent API calls.
3.3.7 Architectural Rationale and Modeling Principles
Discussion: Modeling of State Transitions
Our framework is designed to handle various types of legislative state transitions with precision. A notable example is the handling of revocations. A recommended practice for maintaining a continuous version timeline is to model a revocation not just as the termination of the last valid Version, but also as the creation of a new, subsequent Version with a specific textual content, such as "(Revoked)". However, the API is flexible enough to model this differently. This flexibility is fully represented in the Action data model, which can cleanly represent an original creation (with only a produces_version_id), a simple repeal (with only a terminates_version_id), or a standard amendment (using both).
Discussion: Atomic Actions and Emergent Event Cohesion
A core design principle of the SAT-Graph is the atomicity of causal events. Our model defines an Action as the most granular unit of change possible to ensure maximum auditability. A key consequence is that a single, complex legislative event—such as the enactment of "Constitutional Amendment No. 99"—is not represented by a monolithic node. Instead, its total effect is captured by an emergent collection of many individual Action nodes, whose cohesion is established by their respective source_version_id properties all tracing back to the same source Work.
This design provides the best of both worlds. An agent can use a single, high-level API action to reconstruct the full scope of an event:
  • Strategy: Aggregate Actions by Source Work. The agent first identifies the amending law as a Work (e.g., via resolveItemReference). It then makes a single call to getActionsBySource( source_work_id=...) to gather all the individual Action nodes authorized by that amendment.
This powerful aggregation enables high-level impact analysis ("What were all the effects of this amendment?"), while the atomic nature of each returned Action preserves the granularity needed for pinpoint-accurate causal tracing.
Discussion: Multi-Aspect Retrieval via TextUnit
The TextUnit is the foundational object for the multi-aspect retrieval capability of the SAT-Graph. By decoupling textual content from the core graph entities, we can associate multiple, distinct textual representations with any node. This is achieved via the source_node_type and source_node_id properties on the TextUnit, allowing it to point to any other entity in the graph. For example, an entire Work can have a TextUnit with an aspect of "textual_metadata", a Theme can have one with an aspect of "description", and a Version can have both a "canonical" text and a "summary".
This flexibility is directly exposed through the API, particularly via the aspects parameter in functions like searchTextUnits. All TextUnits, regardless of their source entity or aspect, are candidates for embedding and semantic retrieval. This creates a rich, multi-faceted search space where an agent can retrieve information based not just on the law’s formal content, but also on its context, metadata, and conceptual definitions in a single, unified query.
3.4 Practical Implementation Considerations
While the actions specified above define the conceptual and logical interface to the SAT-Graph, a production-grade implementation should include additional features for efficiency and robust operation.
Batch Operations
To mitigate network latency in scenarios requiring numerous lookups—such as reconstructing the state of a constitutional chapter with 50 articles at a specific date, which would otherwise require 50 sequential API calls—, such as resolving the point-in-time structure of a complex item or analyzing its full history, the API should provide batch-enabled versions of its core retrieval actions. Key batch actions include:
  1. 25.
    getBatchItems(ids: list[ID]) $\rightarrow$ list[Item]
    • Description: A batch-optimized version of ‘getItem‘. Given a list of ids, it efficiently retrieves the full data object for each corresponding Item in a single request.
    • Parameters:
      • ids: A list of ids for the items of interest.
    • Returns: A list of the requested Item objects. The returned list may be shorter than the input list if some ids were not found.
    • Discussion: This is a fundamental utility action used after any discovery process. When an agent obtains a list of ids from functions like ‘searchItems‘ or ‘enumerateItems‘, this batch action allows it to hydrate those references into full objects to access their labels, types, or hierarchical relationships without incurring the high cost of N+1 individual requests.
  2. 26.
    getBatchValidVersions(item_ids: list[ID], timestamp: date-time,
    policy: TemporalPolicy) $\rightarrow$ list[Version]
    • Description: A batch-optimized version of ‘getValidVersion‘. Given a list of item ids, it efficiently retrieves all of their corresponding Versions that were valid on a single, specified date.
    • Parameters:
      • item_ids: A list of ids for the items of interest.
      • timestamp: The specific date for which to check validity for all items.
      • policy: The temporal resolution policy to apply uniformly.
    • Returns: A list of the valid Version objects. Note that the returned list may be shorter than the input list if some items did not have a valid version on the given date.
    • Discussion: This action is essential for efficiently reconstructing the state of a complex hierarchical item at a specific point in time. It is designed to be used directly with the output of ‘enumerateItems‘, avoiding the "N+1 query problem" and drastically improving performance.
  3. 27.
    getBatchTexts(version_ids: list[ID], language: string,
    aspects?: list[string]) $\rightarrow$ list[TextUnit]
    • Description: A batch-optimized version of getTextForVersion. Given a list of Version identifiers, it efficiently retrieves their corresponding textual content.
    • Parameters:
      • version_ids: A list of identifiers for the Versions of interest.
      • language: The language to retrieve.
      • aspects: (Optional) A list of textual aspects to retrieve for all Versions (e.g., [‘"canonical"‘, ‘"summary"‘]). Defaults to [‘"canonical"‘].
    • Returns: A list of the requested TextUnit objects.
    • Discussion: This action is critical for use cases that require inspecting the content of multiple versions, such as finding the exact point of a textual change within a long version history. It prevents the severe performance degradation that would result from iterating and calling getTextForVersion individually for each version.
  4. 28.
    getBatchActions(ids: list[ID]) $\rightarrow$ list[Action]
    • Description: A batch-optimized version of ‘getAction‘. Given a list of action identifiers, it efficiently retrieves the full data object for each corresponding Action in a single request.
    • Parameters:
      • ids: A list of unique identifiers for the actions of interest.
    • Returns: A list of the requested Action objects. The returned list may be shorter than the input list if some identifiers were not found.
    • Discussion: This action is essential for the "hydration" step of analytical queries. After an agent discovers a set of relevant ‘Action‘ IDs via a function like ‘summarizeImpact‘, this batch operation allows it to retrieve the full details of each action (such as their specific ‘source_version_id‘) to build a complete and auditable report without the high performance cost of N+1 individual requests.
3.5 API as a Foundation for Agentic Reasoning
The specified actions collectively form a comprehensive yet controlled interface to the legal knowledge graph. They are intentionally designed to be low-level. An agentic framework is not expected to simply call one action to answer a user’s query. Instead, a planning module is responsible for decomposing a complex question into a sequence or a Directed Acyclic Graph (DAG) of these actions. The power of this approach is best illustrated with examples.
Example 1: Point-in-Time Retrieval Plan
A fundamental question like "What was the text of Article 6 of the Constitution in 1999?" would be decomposed into a clean, sequential plan. The agent’s goal is to deterministically find the single correct textual version for that specific date.
  1. 1.
    Ground the reference: It first calls resolveItemReference(referenceText="Article 6 of the Constitution") to get the item’s canonical ID.
  2. 2.
    Find the valid version: It then calls getValidVersion, passing the item’s ID and the target timestamp ("1999-01-01T12:00:00Z"), to get the unique historical Version object.
  3. 3.
    Retrieve the text: Finally, it calls getTextForVersion, passing the retrieved version’s ID and the desired language (e.g., "pt-BR"), to get the final, correct TextUnit.
This simple chain demonstrates how a natural language query is translated into a fully deterministic and auditable retrieval plan once the initial probabilistic grounding is complete.
Example 2: Thematic Analysis
A broader, exploratory question like "Summarize the evolution of all constitutional provisions related to ’Digital Security’ since 2000" would require a thematic analysis plan that leverages the API’s powerful server-side aggregation:
  1. 1.
    First, the agent discovers the relevant conceptual node by calling searchThemes(semantic_ query="Digital Security").
  2. 2.
    Then, instead of fetching all items and their individual histories ( which would require O(N) calls where N is the number of articles in the chapter—potentially hundreds of API calls for a single constitutional chapter), the agent makes a single, powerful call to summarizeImpact, passing the theme’s id and the specified time interval.
The server performs the complex aggregation, returning a concise ImpactReport. This report contains the aggregate statistics and the identifiers of all relevant changes. If a more detailed narrative is required, the agent can then make targeted calls to getBatchActions and getBatchItems to hydrate the specific details needed for the final summary. This demonstrates how the API transforms a potentially massive client-side processing task into a single, efficient query.
Example 3: Robustness and Multilingual Fallback
The API’s atomic nature enables agents to build robust and user-friendly behaviors, such as multilingual fallback. An agent serving a user who prefers Portuguese ("pt-BR") can be programmed to handle cases where a specific text is not available in that language. Instead of simply failing, the agent can implement a graceful fallback strategy:
  1. 1.
    The agent first attempts to retrieve the text in the user’s preferred language using getTextFor Version(version_id, "pt-BR").
  2. 2.
    If this call fails (e.g., by returning a 404 error), the agent does not give up. It makes a second call to retrieve the text in a default fallback language, such as English: getTextForVersion(version_id, "en").
  3. 3.
    When presenting the result, the agent can now inform the user about the substitution: "The requested text was not available in Portuguese. Displaying the English version instead."
This pattern, enabled by the API’s clear separation of concerns, demonstrates how composable primitives can be used to build sophisticated, resilient, and transparent agentic behaviors that go beyond simple query-response loops.
3.6 Formal OpenAPI Specification
The complete, machine-readable specification of this API is provided as an OpenAPI 3.0 YAML document, available in the project repository [^1] https://github.com/hmartim/sat-graph-api. The OpenAPI specification includes: Complete request/response schemas for all actions; Example requests and responses; Detailed error response specifications; Authentication requirements; Rate limiting policies.
Developers can use this specification to generate client SDKs, test suites, and interactive documentation using standard OpenAPI tooling.
4 Use Cases and Application
To demonstrate the practical utility and expressive power of the Canonical Action API, this section presents three representative use cases. These scenarios are designed to highlight how an agentic framework can leverage the API to answer complex legal questions that are intractable for standard RAG systems. Each use case follows a pattern: a user’s query is presented, the challenge for conventional systems is outlined, and an executable plan composed of our API actions is detailed.
4.1 Use Case 1: Point-in-Time Comparison and Causal Pinpointing
User Query: "What were the exact textual differences in Article 6 of the Brazilian Constitution before and after the amendment that introduced the right to ’housing’ (direito à moradia)?"
Challenge
A standard RAG system would likely retrieve multiple, conflicting versions of Article 6 from its index, unable to deterministically isolate the versions immediately preceding and succeeding a specific conceptual change. It lacks the causal understanding to link the introduction of a word to a specific legislative event.
Agent Execution Plan
The agent’s plan is to find the specific Action that introduced the change, which inherently contains the identifiers for the "before" and "after" states.
  1. 1.
    Identify the Item: The agent first grounds the query by resolving the textual reference to its canonical identifier.
    candidates = resolveItemReference(reference_text="Article 6 of the Constitution")
    target_item_id = candidates[0].id // Agent proceeds with the top candidate
  2. 2.
    Retrieve the Full Historical Lineage: The agent fetches the complete, time-ordered history of all legislative events that affected the item.
    history = getItemHistory(item_id=target_item_id) // Returns a list[Action]
  3. 3.
    Pinpoint the Pivotal Action (Agent Logic): The agent must now find the first Action in the history whose resulting text contains the keyword "moradia". To do this efficiently, it first collects all produces_version_ids from the ‘history‘ list. It then makes a single getBatchTexts call to retrieve all corresponding texts at once. By iterating through these texts, it can identify the pivotal Action responsible for the change. By iterating through the returned texts and their associated source_node_ids, the agent can map the first occurrence of the keyword back to the pivotal Action.
  4. 4.
    Compare the "Before" and "After" States: The pivotal Action object itself contains the direct references to the state it terminated and the state it created. The agent uses these to get a granular difference report.
    // pivotal_action is the Action object found in the previous step
    version_id_before = pivotal_action.terminates_version_id
    version_id_after = pivotal_action.produces_version_id
    diff_report = compareVersions(
    version_id_a=version_id_before,
    version_id_b=version_id_after
Synthesized Outcome
The agent receives a structured TextDiffReport. This verifiable data, along with details from the pivotal_action itself (like its date and source_version_id), is passed to an LLM. The model can then generate a precise, auditable, and natural-language summary: "The right to ’housing’ (moradia) was introduced into Article 6 by an action on [date from action], which terminated the validity of version [ID before] and created version [ID after]. The exact change was the addition of the word ’moradia’…". This answer is precise, verifiable, and causally grounded, demonstrating a capability far beyond standard RAG systems.
4.2 Use Case 2: Causal Lineage Tracing for Legal Audit
User Query: "A legal case from 2012 references Article 227, paragraph 4 of the Constitution. Which specific law introduced this paragraph, and is its text still the same today?"
Challenge
This query has two distinct requirements that are impossible for standard systems: determining historical provenance ("which law introduced…") and confirming current validity with precision. A standard RAG system lacks a causal model to trace origins and cannot guarantee that a retrieved text is the most current, valid version.
Agent Execution Plan
The agent recognizes two parallelizable sub-tasks and starts by grounding the reference to get a stable identifier.
// Initial, common step for both tasks
candidates = resolveItemReference(
reference_text="Article 227, paragraph 4 of the Constitution"
target_item_id = candidates[0].id // Agent proceeds with the top candidate
The agent can now execute the two tasks in parallel.
Task A: Determine Provenance
  1. 1.
    Retrieve the Past Version: The agent gets the specific Version that was valid on a given date in 2012.
    // Timestamp is in ISO 8601 UTC format for consistency
    version_2012 = getValidVersion(
    item_id=target_item_id, timestamp="2012-01-01T12:00:00Z"
  2. 2.
    Trace Causality to the Source: It then traces this specific version back to the action that created it.
    causality = traceCausality(version_id=version_2012.id)
    creating_action = causality.creating_action
  3. 3.
    Hydrate the Source for Readability (Agent Logic): The creating_action object contains the source_version_id. To provide a human-readable answer, the agent "hydrates" this ID (via getVersion and then getItem) and traverses its ancestors (via getAncestors) to find the root Work (the amending law) and get its label.
Task B: Determine Current State
  1. 1.
    Retrieve the Current Version: Using the same item ID, the agent fetches the version valid as of "now".
    // "NOW" is represented by the current UTC timestamp
    version_current = getValidVersion(
    item_id=target_item_id, timestamp="2025-10-06T10:30:00Z"
  2. 2.
    Compare Historical and Current Versions: To check if the text has changed, the agent compares the version from 2012 with the current one.
    diff_report = compareVersions(
    version_id_a=version_2012.id,
    version_id_b=version_current.id
    // If the diff_report is empty, the text is unchanged.
Synthesized Outcome
The agent’s planner combines the results. From Task A, it has the identity of the amending law. From Task B, it knows whether the text has been altered. This structured data is passed to an LLM, which can now generate a highly accurate and useful answer: "The version of Article 227, paragraph 4, valid in 2012, was introduced by **[Name of Amending Law from Task A]**, effective on [date from creating_action]. A comparison with the current version shows that the text **has not changed**. It reads: ’…’". This demonstrates how an agent can compose API primitives to build a complete and auditable answer to a complex, multi-part question.
4.3 Use Case 3: Hierarchical Impact Summarization
User Query: "Provide a summary of all legislative changes, specifically creations and revocations, within the ’National Tax System’ chapter of the Constitution between January 1, 2019, and December 31, 2022."
Challenge
This query is impossible for systems that are not hierarchy-aware. A flat-text system would perform a keyword search for "tax," retrieving irrelevant articles and missing the implicit scope of the "National Tax System" chapter. Furthermore, it cannot differentiate between different types of legislative changes.
Agent Execution Plan
The agent leverages the API’s powerful server-side aggregation capabilities to construct a highly specific and efficient query plan.
  1. 1.
    Identify the Scope: The agent first resolves the reference to the chapter to get its canonical id.
    candidates = resolveItemReference(
    reference_text="National Tax System chapter of the Constitution"
    scope_id = candidates[0].id
  2. 2.
    Request a Filtered Impact Summary: The agent makes a single call to the aggregate analysis action. It constructs a request object and sends it as the body of a POST request to the summarizeImpact endpoint.
    // The agent constructs a request body for a POST call
    request_body = {
    "item_ids": [scope_id],
    "time_interval": [
    "2019-01-01T00:00:00Z",
    "2022-12-31T23:59:59Z"
    "action_types": ["Creation", "Revocation"]
    report = summarizeImpact(body=request_body)
  3. 3.
    Hydrate Details for Reporting (Agent Logic): The returned ImpactReport contains lightweight lists of identifiers. To build a rich, human-readable narrative, the agent "hydrates" these IDs into full objects using efficient batch calls.
    // The response report contains direct lists of IDs
    action_ids = report.actions
    affected_item_ids = report.affected_items
    // Get full data for the legislative events
    full_actions = getBatchActions(ids=action_ids)
    // Get full data for the affected constitutional articles
    full_items = getBatchItems(ids=affected_item_ids)
    // Additional hydration can be done to find the names of the amending laws
Synthesized Outcome
The agent now possesses a complete, structured dataset of all relevant changes. This verifiable data is passed to an LLM for synthesis. The resulting summary is not a guess, but a factual report grounded in the graph: "Between 2019 and 2022, the ’National Tax System’ chapter saw the following changes matching your criteria: Article 156-A was created by Constitutional Amendment No. 110/2019, and **Article 149-A** was revoked by Constitutional Amendment No. 109/2021." This demonstrates a complete, efficient, and auditable workflow, from high-level aggregation to detailed, human-readable reporting.
4.4 Use Case 4: Distinguishing Structural and Normative Predecessors
User Query: "What was the previous rule that governed this matter?"
Challenge
This seemingly simple question hides a deep legal ambiguity. "Previous rule" can have two distinct meanings: (1) the structural predecessor (the immediately preceding version of the same article) or (2) the normative predecessor (an entirely different article, possibly in another law, that was revoked and replaced). Standard systems can handle neither, as they lack both versioning and a semantic understanding of legislative succession.
Agent Execution Plan
Our API empowers an agent to disambiguate and solve for both interpretations. After establishing the context (getting the ‘target_item_id‘ for "this matter"), it can pursue both paths.
For the Structural Predecessor (Deterministic Causal Path):
The agent’s plan is to find the direct causal antecedent of the current version by tracing the event that created it.
  1. 1.
    It retrieves the current version of the item:
    current_version = getValidVersion(item_id=target_item_id,
    timestamp="2025-10-06T10:30:00Z", policy="SnapshotLast")
  2. 2.
    It then traces this version’s immediate cause by calling traceCausality:
    causality = traceCausality(version_id=current_version.id)
  3. 3.
    The returned CausalityTrace object contains the Action that created the current version. This Action also holds the ID of the version that it terminated. This is the structural predecessor:
    predecessor_version_id = causality.creating_action.
    terminates_version_id
This plan is fully deterministic and provides a precise, auditable answer by directly traversing the causal graph primitives.
For the Normative Predecessor (Heuristic Path):
The agent executes a more complex, discovery-oriented plan to find a semantically similar but structurally distinct predecessor.
  1. 1.
    It first gathers information about the current version: its text, and the start of its validity.
    current_version = getValidVersion(
    item_id=target_item_id, timestamp="2025-10-06T10:30:00Z", policy="SnapshotLast"
    change_date = current_version.validity_interval[0] // Start of the interval
    current_text = getTextForVersion(version_id=current_version.id, language="pt-BR").content
  2. 2.
    It then uses this information to search for semantically similar texts that were valid immediately before the change occurred.
    // Agent constructs a request body for a POST call
    // It calculates a timestamp immediately prior to the change_date
    previous_timestamp = calculate_prior_time(change_date) // e.g., change_date - 1 second
    request_body = {
    "semantic_query": current_text,
    "timestamp": previous_timestamp
    candidates = searchTextUnits(body=request_body)
  3. 3.
    The agent can then analyze the top candidates, filtering out the structural predecessor (which might also appear) and presenting the most likely normative predecessor(s) to the user.
This second plan, while heuristic, is powerfully grounded in the verifiable temporal snapshots provided by the API, making its discovery process far more reliable than a blind search.
Synthesized Outcome
The agent can now provide a sophisticated, multi-faceted answer that addresses the user’s ambiguous query head-on: "The query ’previous rule’ can have two meanings. The direct previous version of **Article 10** (its structural predecessor) had the following text: ’…’. However, this new version was created by Law XYZ, which also revoked **Article 150 of a different law** that previously governed this matter (its likely normative predecessor). Would you like to see the text of Article 150?". This demonstrates the framework’s ability to support the nuanced reasoning required for deep legal analysis.
5 Conclusion
In this paper, we addressed a critical challenge in the application of AI to the legal domain: the need for a trustworthy and auditable method for interacting with complex, versioned knowledge. While our previous work, the SAT-Graph RAG, established a robust model for representing the structural and temporal reality of legal norms, it left open the question of how to query this representation without sacrificing its inherent guarantees. We have demonstrated that the solution is not a more complex, monolithic query engine, but rather a more disciplined and transparent interaction layer.
Our primary contribution is the formal specification of the SAT-Graph API. The core of this API is a library of canonical actions—low-level, composable, and deterministic primitives that enforce auditability by design. By providing this API as the primary, formally-specified interface to the graph, we create a powerful abstraction layer. This ensures that agentic frameworks can operate on the knowledge base safely and predictably, without needing to understand the complexities of the underlying data model, making the retrieval process as rigorous as the knowledge base itself. We have shown how this architecture enables a powerful synergy: advanced agentic frameworks can provide the flexibility to interpret complex natural language queries, while our API provides the unyielding determinism required to execute the resulting retrieval plans safely. The agent’s plan becomes a complete, human-readable audit trail, transforming a potentially opaque reasoning process into a series of explicit, verifiable steps. This shift from probabilistic retrieval to deterministic, planned execution is a direct architectural response to the challenge of Explainable AI (XAI), making it possible to build trustworthy AI systems for regulated domains.
5.1 Implications
The implications of this two-layer architecture are significant, marking a paradigm shift away from end-to-end, black-box systems and toward a more modular and trustworthy approach.
Decoupling Knowledge from Reasoning
The most crucial implication is the decoupling of the knowledge representation from the reasoning logic. The SAT-Graph serves as the stable, verifiable "source of truth," which can be curated and maintained independently. The reasoning layer—the agentic frameworks that plan and execute queries—can evolve and improve at a rapid pace, incorporating newer and more powerful LLMs.
Generalizability Beyond the Legal Domain
Furthermore, our deliberate use of a generic Item data model—representing not just a document (‘Work‘) and its components (‘Work Component‘), but also abstract Concepts and Terms—means that the SAT-Graph and its accompanying API are not limited to the legal domain. This architectural pattern is directly applicable to other fields that rely on versioned, structured, and evolving documents, such as technical standards (ISO, IETF), regulatory compliance documents (FDA, financial regulations), and complex engineering manuals. In essence, any domain where what was is as important as what is can benefit from this architectural pattern.
Architectural Boundaries
Finally, it is critical to distinguish between the responsibilities of the SAT-Graph API and those of the higher-level agentic framework that consumes it. Our API is intentionally designed as a set of verifiable primitives for retrieving structural, temporal, and causal facts, answering "what is?" and "what was?". Complex legal reasoning tasks, such as conflict detection or transitive dependency analysis, are the responsibility of the reasoning agent. This deliberate separation of concerns is a core strength of our architecture: the API provides the stable and auditable foundation of truth, while the agent provides the flexible, evolving layer of strategic reasoning.
5.2 Limitations
While this work provides a formal specification and demonstrates the conceptual power of the SAT-Graph API through use cases, several limitations should be acknowledged:
Implementation Complexity
: Building a production-grade implementation requires significant engineering effort, particularly for the server-side aggregation and temporal indexing capabilities.
Ontology Dependency
: The API’s effectiveness depends on the quality and completeness of the underlying SAT-Graph. Incomplete or incorrectly modeled events will propagate to retrieval results.
Agent Planning Capability
: The architecture assumes the availability of sophisticated planning agents. Current LLMs vary widely in their ability to decompose complex queries into optimal action sequences.
Evaluation Gap
: This paper focuses on specification and design. Comprehensive empirical evaluation against legal AI benchmarks remains future work.
5.3 Future Directions
This work lays a foundational layer, and in doing so, opens several avenues for future research.
First, the API specified in this paper focuses on the "verifiable" graph of hierarchical, temporal, and causal links. As highlighted by use cases like impact analysis, a critical next step is the formalization of a parallel "semantic overlay" graph. The conceptual groundwork for this involves introducing a first-class Relation entity, capable of modeling a rich vocabulary of typed, directed, and potentially time-bound links between Items (e.g., ‘cites‘, ‘succeeds‘, ‘applies_to‘). This would be exposed through a new set of ‘Graph Traversal Actions‘ in the API, enabling the tracing of transitive impacts and other complex relational queries. The practical design and formal OpenAPI specification for the Relation entity and its associated actions are being actively developed as part of an open-source project, which serves as the reference implementation for the concepts discussed in this paper. [^2] The SAT-Graph API specification project is publicly available at: https://github.com/hmartim/sat-graph-api
Second, the sophistication of the planner agent is a key area for improvement. Future work could explore fine-tuning LLMs specifically for the task of translating complex questions into optimal execution DAGs using our API. This could involve creating a synthetic dataset of questions and their corresponding gold-standard action plans for various structured document domains.
Finally, this framework can be extended to handle cross-document analysis across an entire corpus. This would involve connecting multiple SAT-Graphs and introducing actions that can trace references and dependencies not just within a single document, but between different standards, regulations, and their historical precedents, moving closer to a comprehensive model of a complete knowledge ecosystem.