### v2

# Deterministic Legal Agents: A Canonical Primitive API for Auditable Reasoning over Temporal Knowledge Graphs

Abstract
For autonomous legal agents to operate safely in high-stakes domains, they require a foundation of absolute determinism and auditability—guarantees that standard Retrieval-Augmented Generation (RAG) frameworks cannot provide. When interacting with temporal knowledge graphs that model the complex evolution of legal norms, agents must navigate versioning, causality, and hierarchical structures with precision, a task for which black-box vector search is ill-suited. This paper introduces a new architectural pattern to solve this: a formal Primitive API designed as a secure execution layer for reasoning over such graphs. Instead of a monolithic query engine, our framework provides a library of canonical primitives—atomic, composable, and auditable primitives. This design empowers planner-guided agents to decompose complex legal questions into transparent execution plans, enabling critical tasks with full verifiability, including: (i) precise point-in-time version retrieval, (ii) robust causal lineage tracing, and (iii) context-aware hybrid search. Ultimately, this architecture transforms opaque retrieval into auditable reasoning, turning the agent’s internal process from a black box into a verifiable log of deterministic primitives and providing a blueprint for building the next generation of trustworthy legal AI.
Keywords: Legal AI, Deterministic Legal Agents, Agentic Frameworks, Temporal Knowledge Graphs, SAT-Graph, Auditability, Explainable AI (XAI), Trustworthy AI, Retrieval-Augmented Generation (RAG).
1 Introduction
The emergence of agentic frameworks powered by Large Language Models (LLMs) has catalyzed a new paradigm in automated reasoning, with Retrieval-Augmented Generation (RAG) serving as a dominant architecture for grounding agent outputs in factual data [1]. While RAG helps reduce hallucinations, the direct application of its standard "flat-text" model by agents in high-stakes, structured domains such as law reveals fundamental limitations. Legal corpora are not mere collections of text; they are complex, interlinked systems defined by a rigid hierarchical structure, a continuous temporal evolution, and a clear causal chain of legislative actions. Recent studies confirm that traditional RAG frameworks, which rely on surface-level semantic similarity, are insufficient for building reliable, domain-specific agents without significant adaptation [2, 3].
Standard RAG architectures are ill-equipped to provide the guarantees these agents require. Even when enhanced with sophisticated semantic segmentation, the underlying retrieval mechanism remains fundamentally flawed. It treats well-defined legal segments as a simple "flat-text" collection. More critically, it lacks any inherent understanding of versioning. A query about a legal provision is meaningless without a temporal anchor, as the correct answer changes over time. This structural and causal blindness is not merely a technical limitation; it is a critical vulnerability that renders any agent relying on it fundamentally unreliable for applications demanding legal precision and auditability.
To address these deficiencies, our prior work introduced the Structure-Aware Temporal Graph RAG (SAT-Graph RAG) [4]. This framework moves beyond flat text by modeling legal norms as a verifiable knowledge graph, as illustrated in Figure 1. While the SAT-Graph provides the necessary verifiable knowledge substrate, a critical gap remains: how can an autonomous agent reliably and flexibly interact with this structured knowledge without sacrificing its deterministic properties? Unstructured access could allow the agent to bypass the graph’s guarantees, undermining the very trustworthiness it was designed to provide.
Figure 1: From "flat-text" RAG to SAT Graph RAG
This paper bridges that critical gap. We argue that enabling trustworthy legal agents requires more than a better knowledge base; it demands a formally defined and auditable interaction protocol. To this end, we introduce a formal query execution layer, which we specify as the SAT-Graph API—a canonical primitive interface for querying structure-aware, temporal graphs with deterministic retrieval guarantees. The core of this API is a library of canonical primitives—atomic, composable, and auditable primitives designed to serve as the secure toolset for agentic reasoning. Instead of a monolithic query engine, we provide the essential building blocks that agents can use to construct complex, verifiable retrieval plans. These primitives provide a comprehensive toolkit for: (i) high-precision hybrid search combining semantic, lexical, and structured metadata filters; (ii) robust resolution of textual references to formal identifiers; (iii) deterministic retrieval of point-in-time versions; and (iv) complete, auditable tracing of causal lineage [^1] It is important to clarify the scope of ”causality” within this framework. We address the chain of legislative events that determine the textual validity of a legal provision at a specific point in time (causality in the textual-documentary world). This is distinct from the legal-dogmatic concept of causality, which typically refers to the relationship between a real-world event (the operative fact) and the legal norm that applies to it..
We demonstrate how a planner-guided agent, such as those proposed in advanced Graph RAG frameworks [5], can leverage this API by decomposing a user’s natural language query into an executable Directed Acyclic Graph (DAG) of these primitives. This two-layer architecture—utilizing a temporal knowledge graph like the SAT-Graph as the verifiable knowledge base, and our API as the secure interaction protocol—establishes a new blueprint for building deterministic legal agents. It is a direct architectural response to the broader challenge of Explainable AI (XAI), moving the agent’s reasoning from an opaque "black box" to a transparent process that can be audited and understood, a critical requirement for high-stakes domains [6]. This allows advanced agents to operate on legal knowledge while ensuring every step of the retrieval process remains explicit, verifiable, and grounded in the structure of the law.
The main contributions of this paper are as follows:
  • We formally specify a Canonical Primitive API, a library of deterministic primitives that serve as the necessary toolset for building auditable agents.
  • We demonstrate how this API serves as a foundational layer for agentic architectures, enabling them to decompose complex questions into verifiable retrieval plans.
  • We introduce a conceptual shift for RAG, transforming the retrieval process from a passive, single-step context injection into an active, agent-driven inquiry where the agent strategically plans and executes the construction of its own context.
  • We articulate how this architecture decouples the knowledge representation (the graph) from the reasoning logic (the agent), enabling a robust and explainable approach to answering legal questions with LLMs.
The remainder of this paper is structured as follows. Section 2 discusses related work in legal AI and graph-based retrieval. Section 3 details the specification of our API. Section 4 presents use cases demonstrating its application, and Section 5 concludes with a discussion of the implications and future directions.
2 Related Work
Our work is situated at the intersection of legal informatics, retrieval-augmented generation, knowledge graphs, and agentic AI systems. We position our contribution by reviewing the state of the art in these areas.
2.1 AI in Legal Information Retrieval
The application of artificial intelligence to legal text is a long-standing field of research. Early systems relied on handcrafted rules and expert knowledge to perform legal reasoning [7]. While the advent of machine learning brought significant advances through models like Legal-BERT [8], a common thread remains the treatment of legal documents as unstructured text. This perspective often overlooks the deep, formal semantics embedded in the law’s structure.
Our approach, in contrast, aligns with a mature research tradition that seeks to formalize legal knowledge using ontologies. The challenge of managing the "diachronic evolution" of legal norms has long been recognized, with early work proposing ontologies to identify and track legal resources over time [9]. Our use of a formal, LRMoo-based model [10] builds upon this foundation. Furthermore, the explicit modeling of legislative changes via our ‘Action‘ nodes is a direct response to the recognized need for formal languages to describe and manage changes in complex knowledge graphs, a problem addressed by frameworks like the Knowledge Graph Change Language (KGCL) [11]. By adopting a formal, top-down ontological model, we enable more sophisticated, machine-readable applications that respect the inherent structure of the law.
2.2 Retrieval-Augmented Generation and its Limitations
Retrieval-Augmented Generation (RAG) has become the de-facto standard for enabling LLMs to answer questions over specific corpora [1]. In its canonical form, RAG operates by embedding and indexing text chunks, retrieving the most relevant via vector similarity search.
While effective for unstructured text, this "flat-text" RAG paradigm falters in the legal domain. Recent benchmarks confirm that the retrieval stage remains a critical bottleneck for legal RAG systems [12]. Our claims of "context blindness" and "temporal naivety" are not hypothetical; they reflect well-documented failure modes. Studies show that a reliance on surface-level similarity fails to capture the "nuanced relevance" of legal statutes [3] and that standard RAG is insufficient without domain-specific adaptations to handle the complexity of legal queries [2]. The inability of these systems to guarantee the retrieval of the correct historical version of a legal norm is a fatal flaw that renders any agent built upon them fundamentally untrustworthy and unsafe for any serious legal application.
2.3 Knowledge Graphs for Retrieval (Graph RAG)
To overcome the limitations of flat-text retrieval, a growing body of work has focused on Graph RAG [13]. Frameworks like Youtu-GraphRAG have demonstrated the power of this approach [5]. The architectural advantage of a knowledge graph is that it provides a persistent, structured, and connected memory that an external agent can query, enabling more intricate reasoning tasks than are possible with isolated text chunks [14].
Our SAT-Graph RAG [4] aligns with this trend but makes a critical distinction. Unlike generalist approaches that discover a graph "bottom-up" from unstructured text, the SAT-Graph is constructed "top-down," mapping the pre-existing, formal structure of the law onto a deterministic, temporal ontology. The most important entities are not named persons or places, but the structural components of the law itself and the legislative events that alter them. While this provides a robust knowledge base, the critical question of how an agent can interact with this knowledge deterministically and auditably remains unaddressed. Without a formal protocol, the guarantees of the graph are easily bypassed, reintroducing the very unreliability we seek to eliminate.
2.4 Agentic Frameworks for Reasoning and Retrieval
Recent advances have seen LLMs employed not just as generators, but as reasoning agents capable of planning and tool use [15]. However, these generic agentic architectures are only as reliable as the tools they are given. In a high-stakes domain like law, a non-deterministic or opaque retrieval tool is a critical point of failure for the entire agentic system. The API we propose is designed to be this missing high-fidelity tool library, providing the deterministic guarantees required for auditable reasoning.
A critical insight from recent research is the symbiotic relationship between architectural innovation and evaluation benchmarks. The legal domain has historically suffered from a lack of benchmarks that require deep, multi-hop reasoning over structured, temporal data. However, new, more challenging benchmarks are emerging, such as "Bar Exam QA" and "Housing Statute QA," which are specifically designed to require "reasoning-intensive" queries rather than simple lexical overlap [16]. These new benchmarks expose the deep limitations of traditional RAG and create a clear need for architectures like ours. The SAT-Graph, combined with our Canonical Primitive API, is precisely the type of verifiable, structure-aware framework needed to successfully address these next-generation evaluation tasks.
2.5 Positioning Our Contribution
Our work synthesizes these threads. We build upon the foundation of our SAT-Graph, which addresses the representation problem in legal AI. We acknowledge the power of agentic reasoning frameworks for flexible query resolution. However, we identify a critical missing component: the secure and auditable bridge between the reasoning agent and the structured knowledge.
While prior work has focused on either the graph structure itself or the high-level agentic logic, our contribution is the formal specification of the intermediate layer: a Canonical Primitive API. This is not another monolithic retrieval system, but a low-level protocol that provides the necessary safety guarantees for the legal domain. It enables advanced agents to perform complex reasoning by transforming retrieval from a single, probabilistic primitive into a deterministic, auditable plan. Every interaction with the knowledge base becomes explicit, verifiable, and grounded in the temporal and structural reality of the law, providing a true foundation for trustworthy legal agents.
3 Specification of the Canonical Primitive API
The central thesis of our work is that for legal agents to be trustworthy, their interaction with a structured knowledge base must be governed by a formal, auditable protocol. We introduce this protocol as the SAT-Graph API, a library of atomic and composable primitives that serve as the agent’s verifiable toolset. This API serves as an abstraction layer, hiding the underlying graph database implementation (e.g., Cypher queries over a Neo4j database). Crucially, it provides a unified interface for both probabilistic discovery primitives (like resolving natural language references) and fully deterministic retrieval primitives that operate on formal identifiers. In this pattern, an autonomous agent is empowered to interact with the knowledge base exclusively through these primitives, ensuring that every step in its reasoning plan is explicit, verifiable, and its level of certainty is understood, as shown in Figure 2.
Figure 2: Diagram illustrating a reasoning agent decomposing a user prompt into tasks that are executed through primitives provided by the SAT-Graph API.
3.1 Design Principles
The design of the API is guided by three core principles that, together, ensure the agent’s trustworthiness, even when its core logic is a probabilistic language model.
  • Maximal Determinism: We acknowledge that any system interpreting natural language has inherently probabilistic components. Our design philosophy is to isolate this non-determinism at the API’s entry points. Primitives for discovery and search are probabilistic by nature. However, the core architectural contribution of this API is to enable an agent to orchestrate fully deterministic reasoning plans once a formal identifier is obtained. This capability for achieving determinism is what distinguishes our approach and justifies the "Deterministic Legal Agent" paradigm. Once a formal id (URI/URN) is acquired, all subsequent primitives that operate on it are guaranteed to be deterministic.
  • Composability: Primitives are designed as atomic "building blocks." They can be chained together by a planning agent to construct a verifiable execution plan. The architectural philosophy is to provide the agent with powerful, deterministic primitives, while the reasoning logic that composes them (e.g., iterating through a list, inspecting text) remains the responsibility of the agent. This deliberate separation ensures that the agent’s entire reasoning process—not just the API calls—is explicit and auditable, preventing opaque server-side operations from obscuring the plan.
  • Verifiability through Auditability: This principle is our safeguard against the initial non-determinism. Every primitive, including the probabilistic ones, must return not just a result but also a confidence score or justification. When an agent executes a plan, the sequence of primitive calls and their outputs forms a complete, human-readable audit trail. This log is the tangible artifact of the agent’s reasoning process, making it fully verifiable, even if the initial discovery steps are probabilistic. A user can inspect the log and see, "The system resolved ’Article 6’ to id ’…’ with 98% confidence," allowing them to trust or question the initial grounding step.
3.2 Core Data Models
The primitives in our API provide an agent with a standardized vocabulary of data objects for perceiving and navigating the knowledge graph. These core objects, which correspond directly to the entities in the SAT-Graph ontology, as illustrated in Figure 3.
Figure 3: SAT-Graph Ontology
Item:
Represents a versioned, structural entity of a legal norm. It is instantiated as one of two concrete subtypes: Work or Work Component.
  • id (ID): A unique, canonical identifier (e.g. URN).
  • type (id): The specific item type id.
  • label (string): The human-readable label.
  • uri (string)?: (Optional) A uniform resource identifier.
  • parent (ID)?: (Optional) The id of its single structural parent. A ‘Work‘ has no parent.
  • children (list[ID])?: (Optional) A list of ids for its structural children. Leaf nodes have no children.
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the item—examples of metadata description of a legal norm’s Work in [17].
Theme:
Represents a conceptual entity used to classify Items and to organize knowledge. Themes can be arranged in their own taxonomy.
  • id (ID): A unique, canonical identifier.
  • label (string): The human-readable label.
  • uri (string)?: (Optional) A uniform resource identifier.
  • parent (list[ID])?: (Optional) A list of ids of broader themes. Themes can have multiple parents. Root themes have an empty list.
  • children (list[ID])?: (Optional) A list of ids of more specific themes.
  • members (list[ID])?: (Optional) A list of ids of Items classified under this theme.
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the theme.
Architectural Note: Separation of Normative and Conceptual Entities
A foundational design principle of this framework is the strict separation of normative entities (Item) from conceptual, classificatory entities (Theme). Items represent the formal, versioned structure of the law itself. Themes, in contrast, form an independent, atemporal ontology used to organize and discover knowledge. An Item’s structural relationships (‘parent‘, ‘children‘) define its fixed position within a document, forming a strict hierarchy. A Theme’s relationships define both its taxonomic position relative to other themes (‘parent‘, ‘children‘) and its classificatory relationship to the law (‘members‘). This clear separation is a critical design choice that empowers the agent. It allows the agent to disambiguate its intent, navigating the formal structure of a norm (its hierarchy) and the conceptual map of the legal domain (its themes) through distinct, unambiguous API primitives.
The ‘Version‘ (Item Temporal Version) and ‘Action‘ data models build upon this DAG foundation to create a fully versioned and causally-linked graph.
While the underlying ontology distinguishes between different semantic types of hierarchy, the API deliberately abstracts these into a conceptually unified pair of properties: parent and children. This pragmatic choice simplifies an agent’s logic, allowing it to reason about "moving up" or "moving down" a hierarchy using the same vocabulary, regardless of the entity type. We acknowledge a minor difference in the data types: an Item has a single parent, while Themes and Versions can have a list of parents. This is an intentional design choice reflecting the semantics of the graph. An Item’s position in a document’s structure is fixed and singular. In contrast, a Theme can belong to multiple broader categories, and a Version can be reused under multiple parent versions as a performance optimization. Despite this type difference, the conceptual abstraction remains consistent: parent always represents a move towards a broader or containing entity, and children a move towards a more specific or contained one. This approach prioritizes a consistent agentic vocabulary without sacrificing the semantic richness of the underlying model.
Version:
Represents a specific, time-bound version of an Item. A Version object is the agent’s key to the past, capturing not only the content but also the structural position of an item at a specific point in time.
  • id (ID): A unique identifier for this specific version.
  • item (ID): The id of the abstract Item it is a version of.
  • validity_interval ([date, date])?: A tuple representing the start and end dates of its validity. The end date is optional (or null) to represent a currently valid or open-ended version.
  • uri (string)?: (Optional) A uniform resource identifier.
  • parent (list[ID])?: (Optional) A list of identifiers for its parent Versions. While a component has one structural parent, this is modeled as a list to support advanced graph optimizations, such as reusing an unchanged child Version under multiple, newly-created parent Versions. A Work’s Version has no parents.
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the version.
Action:
Represents a reified legislative event as the most granular, atomic unit of change in the graph. It is a fundamental building block of the agent’s audit trail, a first-class entity that models a single, verifiable causal link, connecting a specific clause of an authorizing law to the specific change it enacts.
  • id (ID): A unique identifier for the action.
  • type (string): The type of legislative action (e.g., ‘Amendment‘, ‘Revocation‘, ‘Creation‘).
  • date (date): The effective date of the action.
  • source_version (ID): The identifier for the specific Version of the legal text that authorizes this single change.
  • terminates_version? (ID): (Optional) The specific Version whose validity is terminated by this action.
  • produces_version? (ID): (Optional) The specific Version that is created by this action.
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the action.
TextUnit:
Represents a concrete piece of textual information. It is a flexible object used to hold any "aspect" of text associated with any primary entity in the graph. This could be the canonical text of an Version, a detailed description of a Theme, or a textualized metadata aspect of an Item (like a publication info.).
  • id (string): A unique identifier for this specific piece of text.
  • source_node_type (string): The type of the source node (e.g., "Theme", "Item", "Version", "Action"), to allow for unambiguous retrieval by an agent.
  • source_node_id (ID): The unique identifier of the node in the graph to which this text is attached.
  • language (string): The language of the content (e.g., "pt-br").
  • aspect (string): The semantic role of the text relative to its source node (e.g., "canonical" text of a Version, "description" of a Theme, "textual_metadata" of a Work).
  • content (string): The raw textual content.
3.3 Canonical Primitive API Specification
The SAT-Graph API provides an agent with a structured vocabulary of primitives, organized by intent. A core design guarantee governs this vocabulary: any primitive that accepts a formal ID as its primary input is fully deterministic. This clear distinction is critical: it allows the agent to consciously manage uncertainty during its discovery phase, before transitioning to a fully deterministic and auditable execution path. This section details the signature and purpose of each primitive, which serve as the building blocks for the agent’s reasoning plans.
Bridging Baseline RAG with a Temporal Knowledge Graph
A key design challenge is to provide a search interface that is both simple enough to replicate the behavior of a standard, "structurally and temporally-naïve" RAG system, while also exposing the full power of the underlying temporal graph. Traditional RAG pipelines typically index a single, current version of a text. A query like “Which social rights are guaranteed?” executes a semantic search over these current texts.
The SAT-Graph, however, contains all historical versions. A simple semantic search would return a mix of current and outdated, yet relevant, texts. To manage this complexity, the API’s primary search primitive, searchTextUnits, is explicitly designed to give the agent precise control over temporal scope through a simple timestamp parameter:
  • Default (Current State Query): When the timestamp is omitted, the search is automatically constrained to only the currently valid texts, thus replicating the expected behavior of a baseline RAG.
  • Temporal Query: When a timestamp is provided, the search is constrained to the historical snapshot of texts that were valid on that specific date.
This design provides a single, powerful primitive that empowers an agent to seamlessly switch between simple, RAG-like queries and complex historical analysis, depending on its strategic goal.
3.3.1 Discovery and Search Primitives
These primitives serve as the primary entry points for an agent, used when it needs to find entities based on ambiguous natural language references or semantic content.
  1. 1.
    resolveItemReference(reference_text: string, context_id?: ID,
    top_k?: int) $\rightarrow$ list[{ID, confidence: float}]
    • Description: Translates an ambiguous, natural-language reference to a document/component into a ranked list of candidate Item ids. This is the critical first step for any query that names a specific item, acting as the bridge between user language and the graph’s formal identifiers.
    • Parameters:
      • reference_text: A string such as "Article 5 of the Constitution" or "Civil Code".
      • context_id: (Optional) The id of an Item that provides a structural context. This allows the system to accurately resolve relative references and disambiguate ambiguous ones (e.g., resolving "Article 10" to the one within the same law as the context id).
      • top_k: (Optional) The maximum number of candidate matches to return (e.g., defaults to 3).
    • Returns: A ranked list of objects, each containing a probable id and a confidence score indicating the certainty of the match.
    • Discussion: This primitive is indispensable for grounding agentic workflows. Document/component references are often highly ambiguous ("Article 1" could exist in hundreds of laws). This function handles this by returning a ranked list of candidates rather than forcing a single, potentially incorrect, guess. The optional context_id parameter significantly enhances its precision, enabling sophisticated, context-aware interactions. This design is a cornerstone of agent safety. By forcing the agent to explicitly handle a list of candidates and their confidence scores, it prevents catastrophic errors that arise from incorrect initial grounding. This initial, auditable step of uncertainty management is what makes subsequent deterministic plans trustworthy.
  2. 2.
    resolveThemeReference(reference_text: string, top_k?: int) $\rightarrow$
    list[{ID, confidence: float}]
    • Description: Translates a natural-language name or label for a legal theme into a ranked list of candidate Theme ids. This is the primary entry point for grounding a query in a specific conceptual category.
    • Parameters:
      • reference_text: A string representing the name of a theme, such as "Social Security" or "Environmental Law".
      • top_k: (Optional) The maximum number of candidate matches to return (e.g., defaults to 3).
    • Returns: A ranked list of objects, each containing a probable Theme id and a confidence score.
    • Discussion: This primitive is the conceptual counterpart to resolveItemReference. It performs entity linking for thematic nodes, allowing an agent to deterministically identify a theme mentioned by name before using it as a scope in other primitives like searchTextUnits.
  3. 3.
    searchTextUnits(version_ids?: list[ID], item_ids?: list[ID],
    theme_ids?: list[ID], metadata_filter?: MetadataFilter,
    timestamp?: date, semantic_query?: string, lexical_query?: string,
    language?: string, aspects?: list[string], top_k?: int) $\rightarrow$
    list[{text_unit: TextUnit, score: float}]
    • Description: Performs a hybrid search to discover TextUnits nodes by combining semantic, lexical, and structured criteria returning a ranked list of the most relevant results, ready to be used as context for a generation model. This is the primary, optimized entry point for all standard RAG queries.
    • Parameters:
      • version_ids: (Optional) A list of specific Version (Item Temporal Version) identifiers. Provides an explicit, pre-resolved versioned scope. This parameter is mutually exclusive with the combination of item_ids/theme_ids and timestamp.
      • item_ids: (Optional) A list of Item ids. The search is strictly constrained to the provided items. This primitive does not automatically expand to include descendant items; if hierarchical search is desired, the agent is responsible for first constructing the full list of item IDs using a navigation primitive like getItemHierarchy.
      • theme_ids: (Optional) A list of Theme ids. The search is strictly constrained to Items that are direct members of the provided themes. This primitive does not automatically expand to include descendant themes; if hierarchical search is desired, the agent is responsible for first constructing the full list of theme IDs using a navigation primitive like getThemeHierarchy.
      • metadata_filter: (Optional) A structured object (MetadataFilter) that specifies structured filters to be applied. This object can contain clauses like:
          item_metadata_filter (dict): (Optional) A set of key-value pairs to filter against the metadata of the Items (e.g., ‘"jurisdiction": "federal"‘).
          version_metadata_filter (dict): (Optional) A set of key-value pairs to filter against the metadata of the Version itself (e.g., ‘"publication_date": ">=2010-01-01"‘).
      • timestamp: (Optional) An ISO 8601 date-time string to constrain the search. If omitted, the search defaults to the current UTC timestamp at the moment of query execution, effectively retrieving the "presently valid" texts. This parameter is ignored if version_ids is provided.
      • semantic_query (string): (Optional) The natural language query string.
      • lexical_query (string): (Optional) A set of keywords or phrase for a full-text syntactic search.
      • language: (Optional) A code for the desired language.
      • aspects: (Optional) A list of strings specifying the desired textual aspects (e.g., [‘"canonical"‘, ‘"summary"‘]). Defaults to [‘"canonical"‘].
      • top_k: (Optional) The maximum number of results to return.
    • Returns: A ranked list of objects, each containing a matching TextUnit and its similarity score.
    • Discussion: This primitive is unique within the API as it bridges probabilistic discovery and deterministic scoping. While it can accept formal IDs (version_ids), its primary purpose is to perform a search, which is inherently probabilistic. This is an intentional design choice to empower the agent with a powerful tool for focused analysis. This hybrid behavior is critical for sophisticated reasoning. It allows an agent to first build a high-confidence, verifiably correct set of documents (e.g., "all versions of this law during a specific presidential term") and then perform a semantic exploration within that trusted context. This is fundamentally different from a deterministic fetch like getBatchTexts, which would return all texts for the given versions without any semantic ranking or scoring.
  4. 4.
    searchItems(item_ids?: list[ID], theme_ids?: list[ID],
    item_metadata_filter?: dict, semantic_query?: string,
    lexical_query?: string, top_k?: int) $\rightarrow$
    list[{item: Item, score: float}]
    • Description: Performs a hybrid search to discover Item nodes (the structural entities of the graph) by combining semantic, lexical, and structured criteria. This primitive searches across the entire version history of each item’s text to ensure comprehensive discovery.
    • Parameters:
      • item_ids: (Optional) A list of Item ids. The search is strictly constrained to the provided items. This primitive does not automatically expand to include descendant items; if hierarchical search is desired, the agent is responsible for first constructing the full list of item IDs using a navigation primitive like getItemHierarchy.
      • theme_ids: (Optional) A list of Theme ids. The search is strictly constrained to Items that are direct members of the provided themes. This primitive does not automatically expand to include descendant themes; if hierarchical search is desired, the agent is responsible for first constructing the full list of theme IDs using a navigation primitive like getThemeHierarchy.
      • item_metadata_filter (dict): (Optional) A set of key-value pairs to filter against the metadata of the Item (e.g., ‘"jurisdiction": "federal"‘).
      • semantic_query (string): (Optional) The natural language query string.
      • lexical_query (string): (Optional) A set of keywords or phrase for a full-text syntactic search.
      • top_k: (Optional) The maximum number of results to return.
    • Returns: A ranked list of objects, each containing a matching Item and its similarity score.
    • Discussion: This primitive is a critical tool for an agent’s grounding strategy. Its purpose is structural discovery, not content retrieval. When an agent uses searchItems, it is attempting to find the timeless locus of a concept within the graph’s formal structure. This stable, atemporal anchor is the necessary prerequisite for forming deterministic, time-bound queries later on. To support this, the search must be comprehensive, and is therefore intentionally designed to be time-agnostic, considering content from an item’s entire version history.
      The distinction from searchTextUnits is critical:
      • To find the text about "housing" that was valid in 2005, an agent must use searchTextUnits( ..., timestamp="2005-01-01").
      • To find the item that, at any point in its history, has dealt with "housing", an agent must use searchItems.
3.3.2 Deterministic Fetch Primitives
Once an agent has successfully grounded a query to a specific ID, it transitions from probabilistic discovery to a fully deterministic execution path. The following primitives form the core of this path, enabling the agent to retrieve data objects with absolute, non-ambiguous certainty, forming the verifiable steps in its audit trail.
  1. 5.
    getTheme(id: ID) $\rightarrow$ Theme
    • Description: Retrieves the full Theme object given its unique identifier.
    • Parameters:
      • id: The canonical identifier of the theme.
    • Discussion: This primitive exemplifies the API’s core pattern for deterministic, primary-key-based retrieval. For brevity and clarity, we detail only getTheme, but the API provides analogous fetch primitives for all other core data models, including getItem(id: ID), getVersion(id: ID), and getAction(id: ID). These functions are the fundamental tools an agent uses to "hydrate" a referenced ID into its full, verifiable data object, forming a critical step in any reasoning plan.
  2. 6.
    getValidVersion(item_id: ID, timestamp: date-time) $\rightarrow$ Version
    • Description: The core deterministic temporal resolution function. It answers the critical question of ’what was the legislation on this exact date?’ by retrieving the single Version that was legally in force.
    • Parameters:
      • item_id: The id of the item of interest.
      • timestamp: The specific date for which to check validity.
    • Returns: The unique Version object valid on the given date.
  3. 7.
    getTextForVersion(version_id: ID, language: string) $\rightarrow$ TextUnit
    • Description: Deterministically retrieves a specific textual representation (a TextUnit) for a single, known Version. This is a direct-access, primary-key-based lookup, not a search operation.
    • Parameters:
      • version_id: The unique identifier of the Version object whose text is requested.
      • language: A code for the desired language (e.g., "pt-br", "en").
    • Returns: The single, corresponding TextUnit object.
    • Discussion: This primitive is fundamentally different from searchTextUnits. While the latter is a probabilistic discovery tool based on semantic similarity, getTextForVersion is a deterministic fetch operation. It serves the critical use case where an agent, having already identified a specific Version through a deterministic navigation or temporal query, requires its exact textual content. This primitive represents the final step in many reasoning plans: the retrieval of the ground-truth text. Maintaining it as a distinct, primary-key-based primitive is fundamental to the API’s design, ensuring that this final, crucial step is guaranteed to be deterministic and auditable.
3.3.3 Structural Navigation Primitives
These primitives equip an agent with the tools to deterministically navigate the graph’s timeless structures. Unlike search primitives that discover what is relevant, these primitives allow the agent to understand where a piece of information is located and how it relates to its neighbors. This capability is fundamental for constructing context-aware and legally precise reasoning plans.
  1. 8.
    getItemHierarchy(item_id: ID, depth?: int) $\rightarrow$ list[ID]
    • Description: Allows an agent to systematically discover the structural descendants of a known Item. This is the primary deterministic primitive for exploring the graph’s structural hierarchy.
    • Parameters:
      • item_id: The ID of the root Item for the hierarchy traversal.
      • depth: (Optional) An integer controlling the traversal depth. If ‘depth=1‘, only direct children are returned. If omitted or negative, all descendants are returned recursively.
    • Returns: A list of descendant Item IDs.
    • Discussion: This primitive’s purpose is purely structural enumeration. It provides the agent with a deterministic tool for exploration, distinct from probabilistic search. For clarity and brevity, this specification details the hierarchy traversal for Item entities. The same pattern is applied consistently across the API for other hierarchical structures, with analogous primitives such as getVersionHierarchy, getThemeHierarchy and getItemTypeHierarchy available for traversing their respective taxonomies.
  2. 9.
    getItemAncestors(item_id: ID) $\rightarrow$ list[Item]
    • Description: Provides an agent with the full structural lineage of an item, navigating upwards to the root Work. This primitive is fundamental for contextual grounding.
    • Parameters:
      • item_id: The id of the starting Item.
    • Returns: An ordered list of Item objects representing the hierarchical path, from the highest-level ancestor (e.g., a ‘Title‘) down to the immediate parent of the input item. The root Work itself is excluded.
    • Discussion: This primitive is the fundamental tool for contextualization. The legal meaning of a provision is often determined by the Chapter or Title it belongs to. After a search primitive identifies a relevant item, an agent must use getItemAncestors to retrieve its structural ’breadcrumb’. This is not just for generating richer responses; it is a core requirement for explainable reasoning, as the context provided by the ancestors is often part of the answer itself.
  3. 10.
    getThemesForItem(item_id: ID) $\rightarrow$ list[Theme]
    • Description: Retrieves a list of all Themes that are directly associated with a specific Item. This provides the inverse navigation from the structural hierarchy to the conceptual ontology.
    • Parameters:
      • item_id: The id of the Item whose thematic classifications are requested.
    • Returns: A list of all Theme objects that classify the given Item.
    • Discussion: This primitive makes the relationship between items and themes fully bidirectional, enabling an agent to efficiently answer the essential question: ’What are the official conceptual classifications of this law?’ Without this primitive, an agent would need to resort to a highly inefficient brute-force search across the entire thematic ontology. This function completes the API’s structural navigation toolkit.
3.3.4 Causal & Lineage Analysis Primitives
This category of primitives moves beyond static navigation to empower an agent with forensic analysis capabilities. They are the tools for traversing the temporal and causal dimensions of the graph, enabling the agent to deterministically reconstruct the history and evolution of any legal provision.
  1. 11.
    getItemHistory(item_id: ID) $\rightarrow$ list[Action]
    • Description: Provides an agent with the complete, unbroken chain of custody for a legal provision, returning an ordered timeline of every Action that has ever affected it.
    • Parameters:
      • item_id: The id of the Item whose history is requested.
    • Returns: A time-ordered list of all Action objects that have either created or terminated a version (Version) of the specified Item.
    • Discussion: This primitive is the foundation of the agent’s auditability. It provides a complete, verifiable narrative of an item’s evolution. By iterating through the returned list of Action objects, the agent can deterministically reconstruct every past state of the provision, making its historical reasoning transparent and defensible.
  2. 12.
    traceCausality(version_id: ID) $\rightarrow$ {creating_action: Action,
    terminating_action?: Action}
    • Description: Allows an agent to perform pinpoint causal attribution for a single Version. It deterministically identifies the specific Action that created it (its provenance) and, if applicable, the Action that ended its validity.
    • Parameters:
      • version_id: The unique identifier of the Version.
    • Returns: An object containing the Action that created the Version and, optionally, the Action that terminated it.
    • Discussion: For an agent, this is the tool for answering the ’why’ and ’how’ of legal change. It is crucial for pinpointing the exact legislative act responsible for any given state of the law, forming the core of a truly explainable and auditable reasoning process.
  3. 13.
    getVersionsInInterval(item_ids?: list[ID], start_date: date,
    end_date: date) $\rightarrow$ list[Version]
    • Description: Retrieves all Versions of the given Items that were legally valid at any point within a specified time interval.
    • Parameters:
      • item_ids: (Optional) A list of Item ids. The search is strictly constrained to the provided items. This primitive does not automatically expand to include descendant items; if hierarchical search is desired, the agent is responsible for first constructing the full list of item IDs using a navigation primitive like getItemHierarchy.start_date: The start date of the time interval for the query.
      • end_date: The end date of the time interval for the query.
    • Returns: A time-ordered list of all Version objects whose validity period overlaps with the specified interval.
    • Discussion: This function serves as a critical performance optimization for historical analysis. While getItemHistory retrieves the causal events (‘Actions‘), this primitive retrieves the resulting states (‘Versions‘) within a specific temporal window. Replicating this on the client-side would require fetching the entire history and performing complex filtering. By providing this as a server-side primitive, the API enables an agent to efficiently answer questions like "show all versions of this article during a specific presidential term," which is essential for building responsive and scalable legal research applications.
  4. 14.
    compareVersions(version_id_A: ID, version_id_B: ID) $\rightarrow$ TextDiffReport
    • Description: Performs a granular, structure-aware comparison between the textual content of two different Versions of the same Item. It highlights additions, deletions, and modifications.
    • Parameters:
      • version_id_A: Identifier of the first Version.
      • version_id_B: Identifier of the second Version.
    • Returns: A structured TextDiffReport object detailing the changes.
    • Discussion: This primitive provides a powerful, server-side analysis that goes beyond a simple textual ‘diff‘. Because the server has access to the full graph structure, it can perform a structural comparison, identifying not just changed words but also added or removed components (e.g., a new paragraph). This transforms the comparison from a mere text-level operation into a true structural impact analysis. For an agent tasked with explaining a change, this is a critical capability. Centralizing this complex logic on the server provides the agent with a consistent, high-performance tool for generating precise and reliable change summaries.
  5. 15.
    getActionsBySource(source_work_id: ID, action_types?: list[string]) $\rightarrow$ list[Action]
    • Description: An essential primitive for impact analysis that provides forward causal tracing. Given the id of a source Work (e.g., an amending law), it retrieves all atomic Actions authorized by any clause within that Work.
    • Parameters:
      • source_work_id: The id of the source Work (e.g., "Law No. 50 of 2001").
      • action_types: (Optional) A list of strings to filter the results to specific types of actions (e.g., [‘Revocation‘, ‘Creation‘]).
    • Returns: A comprehensive list of all Action objects caused by the specified source that match the filter criteria.
    • Discussion: This primitive is the logical inverse of traceCausality. While the latter traces backwards from a single effect to its granular cause, getActionsBySource traces forwards from a high-level cause (an entire legislative document) to all of its granular effects. It is the primary, high-performance primitive that empowers an agent to answer the critical question of impact analysis: ’What were all the concrete effects of Law X?’. The ability to trace causality forward from a single legislative source to all its distributed effects is a sophisticated reasoning capability now made available as a deterministic primitive.
3.3.5 Introspection & Metadata Primitive
For an agent to be truly autonomous and trustworthy, it must be aware of its environment’s capabilities and limitations. This class of primitives provides this essential capability—a form of programmatic self-awareness. They allow an agent to discover the boundaries of the knowledge graph at runtime, enabling it to formulate more robust plans and interact more intelligently with the user.
  1. 16.
    getTemporalCoverage(item_id: ID) $\rightarrow$ TimeInterval
    • Description: Returns the start and (optional) end dates of the known version history for a specific Item.
    • Returns: A tuple containing the start date of the first Version and the end date of the last known Version. The end date is null if the item is still considered active.
    • Discussion: This primitive enables an agent to be aware of its own knowledge boundaries. It should be used to proactively validate a user’s temporal query before attempting retrieval, allowing for graceful handling of requests for dates outside the known history and improving the overall user experience.
  2. 17.
    getAvailableLanguages() $\rightarrow$ list[string]
    • Description: Returns a list of all language codes (e.g., "pt-br", "en") for which textual content is available in the graph.
    • Discussion: This enables an agent or user interface to dynamically present language options to the user, ensuring that requests are only made for languages that the system actually supports.
  3. 18.
    getSupportedActionTypes() $\rightarrow$ list[string]
    • Description: Returns a list of all action types (e.g., ‘Amendment‘, ‘Revocation‘) recognized by the graph’s ontology.
    • Discussion: This primitive provides the canonical list of values that can be used in the action_ types filter of other functions like getActionsBySource. It allows a client application to dynamically build filtering interfaces without hardcoding ontology-specific values.
  4. 19.
    getRootThemes() $\rightarrow$ list[Theme]
    • Description: Retrieves a list of the full Theme objects that are at the root of the thematic taxonomy (i.e., themes that have no parents).
    • Returns: A list of all root Theme objects.
    • Discussion: This primitive serves as the primary and most efficient entry point for an agent or user interface to explore the thematic ontology. By returning the full root objects, it allows a client to render the top-level categories without requiring subsequent API calls.
3.3.6 Architectural Rationale and Modeling Principles
Discussion: Modeling of State Transitions
Our framework is designed to handle various types of legislative state transitions with precision. This is fully represented in the Action data model, where the optionality of its properties allows an agent to unambiguously distinguish between different event types:
  • A Creation event only has a produces_version_id.
  • A pure Repeal or Revocation event might only have a terminates_version_id.
  • A standard Amendment has both.
This removes any guesswork from the agent’s causal analysis. A recommended modeling practice, for instance, is to represent a revocation not as a pure repeal, but as an amendment that creates a new version with the text "(Revoked)", thus maintaining a continuous version timeline. The API’s flexibility supports both patterns.
Discussion: Atomic Primitives and Emergent Event Cohesion
A core design principle of the SAT-Graph is the atomicity of causal events. Our model defines an Action as the most granular unit of change possible to ensure maximum auditability. A key consequence is that a single, complex legislative event—such as the enactment of "Constitutional Amendment No. 99"—is not represented by a monolithic node. Instead, its total effect is captured by an emergent collection of many individual Action nodes, whose cohesion is established by their respective source_version properties all tracing back to the same source Work.
This design provides the best of both worlds for an intelligent agent. It can use a single, high-level API primitive (getActionsBySource) to efficiently grasp the full scope of a legislative event:
  • Strategy: Aggregate Actions by Source Work. The agent first identifies the amending law as a Work (e.g., via resolveItemReference). It then makes a single call to getActionsBySource( source_work_id=...) to gather all the individual Action nodes authorized by that amendment.
This powerful aggregation empowers the agent to perform high-level impact analysis, while the atomic nature of each returned Action preserves the granular verifiability needed for a trustworthy audit trail. The agent can see both the forest and the trees, deterministically.
Discussion: Multi-Aspect Retrieval via TextUnit
The TextUnit is the foundational object for the multi-aspect retrieval capability of the SAT-Graph. By decoupling textual content from the core graph entities, we can associate multiple, distinct textual representations with any node. This is achieved via the source_node_type and source_node_id properties on the TextUnit, allowing it to point to any other entity in the graph. For example, an entire Work can have a TextUnit with an aspect of "textual_metadata", a Theme can have one with an aspect of "description", and a Version can have both a "canonical" text and a "summary".
This flexibility is directly exposed through the API, particularly via the aspects parameter in functions like searchTextUnits. All TextUnits, regardless of their source entity or aspect, are candidates for embedding and semantic retrieval. This design creates a rich, multi-faceted search space. It empowers an agent with holistic perception, allowing it to discover relevant information based not just on the law’s formal content, but also on its contextual metadata and conceptual definitions in a single, unified query. This prevents the agent from developing ’tunnel vision’, ensuring that its initial, probabilistic discovery phase is as comprehensive as possible before it commits to a deterministic plan.
3.4 Practical Implementation Considerations
While the atomic primitives specified above provide the logical foundation for auditability, a production-grade API must also enable agents to execute complex reasoning plans efficiently. This section details practical considerations, most notably batch operations, which are essential for making sophisticated, multi-step agentic workflows viable in a real-world setting.
Batch Operations
An agent’s reasoning plan is not a single query, but a sequence of primitives. Scenarios like reconstructing the state of an entire constitutional chapter or analyzing a long causal history would require dozens or even hundreds of atomic calls. To make such complex plans executable, the API must provide batch-enabled versions of its core deterministic primitives. These are not mere optimizations; they are fundamental tools that allow an agent to gather large sets of evidence efficiently without flooding the network with N+1 individual requests. Key batch primitives include:
  1. 20.
    getBatchItems(ids: list[ID]) $\rightarrow$ list[Item]
    • Description: A batch-optimized version of ‘getItem‘. Given a list of ids, it efficiently retrieves the full data object for each corresponding Item in a single request.
    • Parameters:
      • ids: A list of ids for the items of interest.
    • Returns: A list of the requested Item objects. The returned list may be shorter than the input list if some ids were not found.
    • Discussion: This is a fundamental utility primitive used after any discovery process. When an agent obtains a list of ids from functions like ‘searchItems‘ or ‘getItemHierarchy‘, this batch primitive allows it to hydrate those references into full objects to access their labels, types, or hierarchical relationships without incurring the high cost of N+1 individual requests.
  2. 21.
    getBatchValidVersions(item_ids: list[ID], timestamp: date-time) $\rightarrow$ list[Version]
    • Description: A batch-optimized version of ‘getValidVersion‘. Given a list of item ids, it efficiently retrieves all of their corresponding Versions that were valid on a single, specified date.
    • Parameters:
      • item_ids: A list of ids for the items of interest.
      • timestamp: The specific date for which to check validity for all items.
    • Returns: A list of the valid Version objects. Note that the returned list may be shorter than the input list if some items did not have a valid version on the given date.
    • Discussion: This primitive is essential for efficiently reconstructing the state of a complex hierarchical item at a specific point in time. It is designed to be used directly with the output of ‘getItemHierarchy‘, avoiding the "N+1 query problem" and drastically improving performance.
  3. 22.
    getBatchTextUnits(requests: list[{source_node_type: string,
    source_node_id: ID, language: string, aspects?: list[string]}])
    $\rightarrow$ list[TextUnit]
    • Description: A highly flexible, batch-optimized primitive that retrieves multiple TextUnits for a variety of source entities in a single request.
    • Parameters:
      • requests: A list of request objects. Each object in the list must specify:
          source_node_type: The type of the source entity (e.g., "Version", "Theme").
          source_node_id: The unique identifier of the source entity.
          language: The desired language code (e.g., "pt-BR").
          aspects: (Optional) A list of textual aspects to retrieve (e.g., ["canonical", "summary"]).
    • Returns: A single list containing all the requested TextUnit objects, preserving the order of the original request list.
    • Discussion: This powerful primitive is critical for agent efficiency. After executing a complex plan and gathering a diverse set of IDs (some Versions, some Themes), an agent can use this single batch call to hydrate all necessary textual context at once, dramatically reducing network latency.
  4. 23.
    getBatchActions(ids: list[ID]) $\rightarrow$ list[Action]
    • Description: A batch-optimized version of ‘getAction‘. Given a list of action identifiers, it efficiently retrieves the full data object for each corresponding Action in a single request.
    • Parameters:
      • ids: A list of unique identifiers for the actions of interest.
    • Returns: A list of the requested Action objects. The returned list may be shorter than the input list if some identifiers were not found.
    • Discussion: This primitive is essential for the "hydration" step of analytical queries. For example, after an agent retrieves all primitives caused by a specific law via getActionsBySource, it might only receive a list of IDs. This batch operation allows it to retrieve the full details of each action (such as their specific terminated and produced versions) to build a complete and auditable impact report without the high performance cost of N+1 individual requests.
3.5 The API as a Grammar for Agentic Reasoning
The specified primitives are not merely a list of endpoints; they are a formal grammar for agentic reasoning. Collectively, they form a comprehensive yet controlled interface to the temporal knowledge graph. By design, they are low-level primitives. An agent is not expected to answer a user’s query with a single primitive. Instead, its planning module must act as a translator, decomposing a complex question into a Directed Acyclic Graph (DAG) of these primitives. The resulting DAG is the agent’s auditable reasoning plan. The power of this architectural pattern is demonstrated in the following section through a series of practical use cases.
3.6 Formal OpenAPI Specification
The complete, optimized and machine-readable specification of this API is provided as an OpenAPI 3.0 YAML document, available in the project repository [^2] Further examples, detailed execution plans, and the full OpenAPI specification are available in the project’s public repository: https://github.com/hmartim/sat-graph-api.. The OpenAPI specification includes: complete request/response schemas for all primitives; detailed examples of context recovery planning etc. Developers can use this specification to generate client SDKs, test suites, and interactive documentation using standard OpenAPI tooling.
4 Use Cases and Application
To demonstrate the practical utility and expressive power of the Canonical Primitive API, this section presents representative use cases. These scenarios are designed to highlight how an agent’s planning module can compose the API’s atomic primitives into verifiable reasoning plans to answer legal questions that are intractable for standard RAG systems. Each use case follows a pattern: a user’s natural language query is presented, and an executable plan composed of our API primitives is detailed.
4.1 Use Case 1: Deterministic Point-in-Time Retrieval
User Query: "What was the text of Article 6 of the Brazilian Constitution on May 20th, 2001?"
Challenge
This seemingly simple question is impossible for a standard, temporally-naive RAG system to answer reliably. A flat-text index would likely contain multiple, conflicting versions of "Article 6" from different points in time. The system lacks the temporal awareness to understand what "on May 20th, 2001" means, and cannot guarantee that the version it retrieves is the one that was legally in force on that specific date. It would likely return the most recent or most semantically similar version, leading to a factually incorrect and misleading answer. [^3] One might argue that a RAG system could be enhanced by indexing each text chunk with temporal metadata (e.g., ‘valid_from‘, ‘valid_to‘) to be used in a pre- or post-retrieval filter. While this improves precision, the core retrieval mechanism remains a probabilistic semantic search over all chunks that satisfy the temporal filter. There would be no guarantee of retrieving the single, canonically correct version, especially if multiple text chunks from different but overlapping sources were valid on the same date. Our approach, in contrast, uses a deterministic, primary-key-based lookup (getValidVersion) on a specific, pre-resolved item, eliminating retrieval uncertainty entirely.
Agent Execution Plan
The agent’s plan is a clean, sequential execution designed to isolate the single, correct version of the text. The entire process becomes deterministic after the initial grounding step.
  1. 1.
    Ground the Reference: The agent first translates the ambiguous natural language reference into a formal, canonical ID. The added specificity "Brazilian Constitution" helps the resolveItem Reference primitive to find the correct entity with high confidence. The agent planner also knows that the canonical text of an article resides in its caput.
    resolveItemReference(
    reference_text="Article 6, caput of the Brazilian Constitution"
  2. 2.
    Find the Valid Version: Using the now-stable ID, the agent makes a deterministic call to getValid Version, passing the item’s ID and the precise target timestamp to retrieve the unique historical Version object that was valid on that day.
    getValidVersion(
    item_id="...",
    timestamp="2001-05-20T12:00:00Z"
  3. 3.
    Retrieve the Ground-Truth Text: Finally, with the unique version ID, the agent retrieves the exact, unaltered text. The context "Brazilian" informs the choice of language. This final step is also fully deterministic.
    getTextForVersion(
    version_id="...",
    language="pt-BR"
Synthesized Outcome
The agent receives a single, verifiable TextUnit object. This structured data—the text itself, along with its version ID and validity interval—is then passed to an LLM. The model can now generate a precise and trustworthy answer: "On May 20th, 2001, the text of Article 6 of the Brazilian Constitution was: ’…’". The answer is not a probabilistic guess but a factual statement, fully grounded in a deterministically retrieved and auditable piece of evidence.
4.2 Use Case 2: Causal Pinpointing and Version Comparison
User Query: "What were the exact textual differences in Article 6 of the Brazilian Constitution before and after the amendment that introduced the right to ’housing’ (direito à moradia)?"
Challenge
This query is impossible for a standard RAG system. It requires not just retrieving text, but understanding causality—linking a conceptual change (the introduction of "housing") to a specific, discrete legislative event. A standard system lacks a causal model; it cannot deterministically find the pivotal amendment, nor can it isolate the precise "before" and "after" states that this event created. It can only search for the keyword, returning a list of contextually jumbled and unverified text chunks.
Agent Execution Plan
The agent’s plan is to perform a historical investigation. It first retrieves the complete lineage of the article and then analyzes it to pinpoint the exact Action responsible for the change. By design, the final inspection of the texts is performed by the agent itself, making this analytical step an explicit part of its auditable reasoning plan, rather than hiding it within an opaque server-side function. This pivotal Action itself contains the direct links to the "before" and "after" versions.
  1. 1.
    Ground the Reference: As in the previous use case, the agent first grounds the query by resolving the textual reference to its canonical identifier.
    resolveItemReference(
    reference_text="Article 6, caput of the Brazilian Constitution"
  2. 2.
    Retrieve the Full Historical Lineage: The agent fetches the complete, time-ordered history of all legislative events that have ever affected the item. This provides the full evidentiary record for its investigation.
    getItemHistory(item_id="...") // Returns a list[Action]
  3. 3.
    Pinpoint the Pivotal Action: The agent must now identify the specific Action that introduced the keyword. To do this efficiently, it avoids N+1 calls by first collecting all produces_versions from the history and making a single getBatchTexts call. It then programmatically inspects the returned texts to find the first occurrence of "moradia" and identifies the corresponding pivotal Action.
  4. 4.
    Compare the "Before" and "After" States: The pivotal Action object itself contains the direct, verifiable references to the state it terminated and the state it created. The agent uses these IDs to request a granular, server-side difference report.
    // pivotal_action is the Action found in the previous step
    compareVersions(
    version_id_a=pivotal_action.terminates_version,
    version_id_b=pivotal_action.produces_version
Synthesized Outcome
The agent receives a structured TextDiffReport. This verifiable data, along with critical metadata from the pivotal_action object itself (such as its effective date and the ID of the provision that authorized it), is passed to an LLM. The model can now generate a precise, auditable, and natural-language summary: "The right to ’housing’ (moradia) was introduced into Article 6 by Article 1 of Constitutional Amendment No. 26, effective on February 14th, 2000. This action [ID before] and [ID after] The exact textual change was the addition of the word ’moradia’ after ’lazer’." This answer is not just correct; it is causally grounded, demonstrating a level of explainability far beyond standard RAG systems.
4.3 Use Case 3: Thematic and Hierarchical Impact Analysis
User Query: "Summarize the evolution of all provisions related to the theme ’Digital Security’ since 2000."
Challenge
This broad, analytical query is completely intractable for systems that are not simultaneously aware of hierarchy, themes, and time. A standard RAG system would likely perform a simple keyword search for "Digital Security," retrieving a mix of relevant and irrelevant provisions while completely missing the implicit scope of a thematic category. It cannot understand that if a high-level provisions is tagged with a theme, all its sub-components are also implicitly part of that scope. The system lacks the multi-faceted awareness to even begin to scope the query correctly, let alone analyze its evolution over time.
Agent Execution Plan
The agent executes a sophisticated, multi-phase plan. It first deterministically resolves the complex, multi-layered scope of the query. Only after this scope is fully defined does it proceed to the historical analysis phase. The plan is designed for maximum efficiency, using parallel execution and optimized, scoped queries.
Phase 1: Deterministic Scope Resolution
  1. 1.
    Ground Thematic Scope: The agent first calls resolveThemeReference to find the canonical ID for the "Digital Security" theme. It then expands this scope by calling a navigation primitive to get all descendant themes.
    resolveThemeReference(reference_text="Digital Security")
    // Followed by getThemeHierarchy(...)
  2. 2.
    Find Anchor Items: The agent performs a single, highly efficient searchItems call that retrieves all items that are linked to the ’Digital Security’ thematic tree. This efficiently finds the high-level "anchor" items.
  3. 3.
    Hierarchical Expansion: For each anchor item found, the agent makes a parallel call to getItem Hierarchy to gather the IDs of all its descendants. This correctly propagates the thematic relevance down the hierarchy (e.g., from a tagged article to all its untagged paragraphs). The union of these sets forms the final, complete list of relevant provision IDs.
Phase 2: Parallel Historical Analysis
  1. 4.
    Fetch Historical Events: Now, with a precise list of hundreds of item IDs, the agent avoids making individual history calls. It strategically groups the IDs by their root document (e.g., all from the 1988 Constitution). For each group, it makes a single, powerful API call to retrieve all Action objects that affected those specific items since the year 2000. These calls are executed in parallel for maximum performance.
  2. 5.
    Aggregate and Synthesize: The agent collects the results from all parallel calls, now possessing a complete, verifiable dataset of every relevant change. This structured data is aggregated to produce statistics (e.g., 5 amendments, 2 revocations) and then passed to an LLM.
Synthesized Outcome
The LLM receives a rich, structured context containing every single piece of evidence. It can then generate a high-level summary grounded in facts, not probabilistic guesses. The response can be both quantitative and narrative: "Since 2000, there have been X legislative actions affecting Y provisions related to ’Digital Security’. The most significant change was Constitutional Amendment No. 115/2022, which elevated data protection to a fundamental right by amending Article 5…" Each claim in this summary can be traced back to a specific Action object retrieved by the agent, making the entire analytical process fully auditable and trustworthy.
5 Conclusion
In this paper, we addressed a critical challenge preventing the safe deployment of autonomous AI in the legal domain: the lack of a trustworthy and auditable method for interacting with complex, versioned knowledge. While knowledge graphs like our SAT-Graph can represent the structural and temporal reality of legal norms, the question of how an agent can query this representation without sacrificing its guarantees remained open. We have demonstrated that the solution is not a more complex, monolithic query engine, but rather a more disciplined and transparent interaction layer: a formal Canonical Primitive API.
Our primary contribution is a new architectural pattern for building deterministic legal agents, specified through this API. The core of the framework is a library of canonical primitives—low-level, composable primitives that enforce auditability by design. This approach creates a powerful synergy: advanced agentic frameworks provide the flexibility to interpret natural language, while our API provides the unyielding determinism required to execute the resulting reasoning plans safely. The agent’s plan itself becomes a complete, human-readable audit trail, transforming its internal reasoning process from an opaque black box into a series of explicit, verifiable steps.
We have shown that this shift—from passive, probabilistic retrieval to active, agent-driven inquiry built on deterministic primitives—is a direct architectural response to the challenge of Explainable AI (XAI). It provides a concrete blueprint for building the next generation of trustworthy AI systems, not just for law, but for any regulated domain where precision, auditability, and historical correctness are paramount.
5.1 Implications
The implications of enforcing a formal Primitive API between an agent and a temporal knowledge graph are significant, marking a paradigm shift away from end-to-end, black-box systems and toward a more modular and trustworthy approach.
Decoupling Knowledge from Reasoning
The most essential implication is the clean decoupling of the knowledge representation (the graph) from the agentic reasoning layer. The SAT-Graph serves as the stable, verifiable "source of truth," which can be curated and maintained independently. The reasoning layer—the agentic frameworks that plan and execute queries—can evolve at a rapid pace, incorporating newer and more powerful LLMs without compromising the foundational integrity of the knowledge base.
Generalizability Beyond the Legal Domain
This Primitive API architecture is not a solution only for law; it is a generalizable pattern for building trustworthy agents in any regulated, structured domain. Our deliberate use of a generic Item data model means that this pattern is directly applicable to other fields that rely on versioned and evolving documents, such as technical standards (ISO, IETF), regulatory compliance (FDA, financial regulations), and complex engineering manuals. In essence, any domain where what was is as important as what is can benefit from this approach to enabling auditable agentic systems.
Architectural Boundaries
Finally, it is critical to distinguish between the responsibilities of the Primitive API and those of the higher-level agentic framework that consumes it. Our API is intentionally designed as a set of verifiable primitives for retrieving structural, temporal, and causal facts. It answers "what is?" and "what was?". Complex reasoning tasks, such as conflict detection or transitive dependency analysis, remain the responsibility of the reasoning agent. This deliberate separation of concerns is a core strength of our architecture: the API guarantees the integrity of the evidence, while the agent is responsible for constructing the argument.
5.2 Limitations
While this work proposes a complete architectural pattern and formal API specification, several limitations and areas for future work should be acknowledged:
Implementation Complexity
Building a production-grade implementation of the server-side logic is a significant engineering effort. The performance of the API, particularly for temporal and causal queries over a large graph, depends on sophisticated indexing strategies that are non-trivial to implement and optimize.
Knowledge Base Dependency
The determinism of the agent’s reasoning is fundamentally bound to the quality and completeness of the underlying knowledge graph. The principle of "garbage in, garbage out" applies: incomplete or incorrectly modeled legislative events in the SAT-Graph will lead to factually correct but legally incomplete answers from the agent. The API provides transactional integrity but cannot create knowledge that is not there.
Agent Planner Sophistication
Our architecture fundamentally shifts the primary challenge from retrieval to agentic planning. The effectiveness of the entire system now relies on the agent’s ability to translate a user’s natural language query into an optimal and correct sequence of primitives. While the capabilities of LLMs as planners are advancing rapidly, generating complex, multi-step, and efficient DAGs of primitives remains a frontier research problem.
Empirical Evaluation Gap
This paper focuses on the formal specification and conceptual power of the architecture. A comprehensive empirical evaluation against emerging legal AI benchmarks that test temporal and causal reasoning is a critical next step. Such an evaluation would be necessary to quantify the performance and accuracy gains of this agentic, primitive-based approach over traditional RAG models.
5.3 Future Directions
By establishing a formal, primitive-based interface for agentic reasoning, this work lays a foundational layer for numerous avenues of future research.
First, we plan to expand the API’s "grammar" to support deeper legal reasoning. The current specification focuses on the verifiable graph of structural, temporal, and causal links. A critical next step is the formalization of a parallel "semantic overlay" graph. This involves introducing a first-class Relation entity to model a rich vocabulary of typed links between Items (e.g., ‘cites‘, ‘succeeds‘, ‘applies_to‘). A new set of ‘Graph Traversal Primitives‘ would then empower an agent to trace transitive impacts and other complex relational queries, moving from "what happened?" to "how are these concepts legally related?". The design for these extensions is being actively developed in our open-source project. [^4] The SAT-Graph API specification project is publicly available at: https://github.com/hmartim/sat-graph-api
Second, as noted in our limitations, the sophistication of the planner agent is now a key determinant of the system’s overall capability. The next challenge is to improve the agent’s fluency in translating complex questions into optimal execution DAGs using our API. Future work will explore fine-tuning LLMs specifically for this "text-to-plan" task. This involves creating a synthetic dataset of natural language questions and their corresponding gold-standard primitive plans, which can be used to teach agents how to reason effectively over structured, temporal domains.